Web2c

for version 7.4.5 January 2003

K. Berry (kb@mail.tug.org)
O. Weber (infovore@xs4all.nl)

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996-2002 K. Berry & O. Weber. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

1 Introduction

This manual corresponds to version 7.4.5 of Web2c, released in January 2003.

Web2c is the name of a TEX implementation, originally for Unix, but now also running under DOS, Amiga, and other operating systems. By TEX implementation, we mean all of the standard programs developed by the Stanford TEX project directed by Donald E. Knuth: Metafont, DVItype, GFtoDVI, BibTEX, Tangle, etc., as well as TEX itself. Other programs are also included: DVIcopy, written by Peter Breitenlohner, MetaPost and its utilities (derived from Metafont), by John Hobby, etc.

General strategy: Web2c works, as its name implies, by translating the WEB source in which TEX is written into C source code. Its output is not self-contained, however; it makes extensive use of many macros and functions in a library (the ‘web2c/lib’ directory in the sources). Therefore, it will not work without change on an arbitrary WEB program.

Availability: All of Web2c is freely available—“free” both in the sense of no cost (free ice cream) and of having the source code to modify and/or redistribute (free speech). (See section “unixtex.ftp” in Kpathsea, for the practical details of how to obtain Web2c.) Different parts of the Web2c distribution have different licensing terms, however, reflecting the different circumstances of their creation; consult each source file for exact details. The main practical implication for redistributors of Web2c is that the executables are covered by the GNU General Public License, and therefore anyone who gets a binary distribution must also get the sources, as explained by the terms of the GPL (see section “Copying” in Kpathsea). The GPL covers the Web2c executables, including tex, because the Free Software Foundation sponsored the initial development of the Kpathsea library that Web2c uses. The basic source files from Stanford, however, have their own copyright terms or are in the public domain, and are not covered by the GPL.

History: Tomas Rokicki originated the TEX-to-C system in 1987, working from the first change files for TEX under Unix, which were done primarily by Howard Trickey and Pavel Curtis. Tim Morgan then took over development and maintenance for a number of years; the name changed to Web-to-C somewhere in there. In 1990, Karl Berry became the maintainer. He made many changes to the original sources, and started using the shorter name Web2c. In 1997, Olaf Weber took over. Dozens of other people have contributed; their names are listed in the ‘ChangeLog’ files.

Other acknowledgements: The University of Massachusetts at Boston (particularly Rick Martin and Bob Morris) has provided computers and ftp access to me for many years. Richard Stallman at the Free Software Foundation employed me while I wrote the original path searching library (for the GNU font utilities). (rms also gave us Emacs, GDB, and GCC, without which I cannot imagine developing Web2c.) And, of course, TEX would not exist in the first place without Donald E. Knuth.

Further reading: See Appendix B [References], page 55.

2 Installation

(A copy of this chapter is in the distribution file ‘web2c/INSTALL’.)

Installing Web2c is mostly the same as installing any other Kpathsea-using program. Therefore, for the basic steps involved, see section “Installation” in Kpathsea. (A copy is

in the file ‘kpathsea/INSTALL’.)

One peculiarity to Web2c is that the source distribution comes in two files: ‘web.tar.gz’ and ‘web2c.tar.gz’. You must retrieve and unpack them both. (We have two because the former archive contains the very large and seldom-changing original WEB source files.) See section “unixtex.ftp” in Kpathsea.

Another peculiarity is the MetaPost program. Although it has been installed previously as mp, as of Web2c 7.0 the installed name is now mpost, to avoid conflict with the mp program that does prettyprinting. This approach was recommended by the MetaPost author, John Hobby. If you as the TEX administrator wish to make it available under its shorter name as well, you will have to set up a link or some such yourself. And of course individual users can do the same.

For solutions to common installation problems and information on how to report a bug, see the file ‘kpathsea/BUGS’ (see section “Bugs” in Kpathsea). See also the Web2c home

page, http://www.tug.org/web2c.

Points worth repeating:

• Before starting the standard compilation and installation you must install the basic fonts, macros, and other library files. See section “Installation” in Kpathsea.

• If you do not wish to use the standard file locations, see section “Changing search paths” in Kpathsea.

• Some Web2c features are enabled or disabled at configure time, as described in the first section below.

2.1 configure options

This section gives pointers to descriptions of the ‘--with’ and ‘--enable’ configure arguments that Web2c accepts. Some are specific to Web2c, others are generic to all Kpathsea-using programs.

For a list of all the options configure accepts, run ‘configure --help’. The generic options are listed first, and the package-specific options come last.

For a description of the generic options (which mainly allow you to specify installation directories) and basic configure usage, see section “Running configure scripts” in Autoconf , a copy is in the file ‘kpathsea/CONFIGURE’.

‘--disable-dump-share’

Do not make fmt/base/mem files sharable across different endian architectures. See Section 3.5.2.3 [Hardware and memory dumps], page 11.

‘--without-maketexmf-default’
‘--without-maketexpk-default’
‘--without-maketextfm-default’
‘--with-maketextex-default’

Enable or disable the dynamic generation programs. See section “mktex configuration” in Kpathsea. The defaults are the inverse of the options, i.e., everything is enabled except mktextex.

‘--enable-auto-core’
          Dump ‘core’ if the input file is ‘HackyInputFileNameForCoreDump.tex’. See

Section 3.5.1.1 [Preloaded executables], page 10.

‘--enable-shared’

Build Kpathsea as a shared library. See section “Shared library” in Kpathsea.

‘--with-editor=cmd’

Change the default editor invoked by the ‘e’ interactive command. See Section 3.5.3 [Editor invocation], page 12.

‘--with-epsfwin’
‘--with-hp2627win’
‘--with-mftalkwin’
‘--with-nextwin’
‘--with-regiswin’
‘--with-suntoolswin’
‘--with-tektronixwin’
‘--with-unitermwin’
‘--with-x’
‘--with-x-toolkit=KIT’
‘--with-x11win’
‘--with-x11’

Define Metafont graphics support; by default, no graphics support is enabled. See Section 5.5 [Online Metafont graphics], page 25.

‘--x-includes=dir’
‘--x-libraries=dir’

Define the locations of the X11 include files and libraries; by default, configure does its best to guess). See section “Optional Features” in Autoconf . A copy

          is in ‘kpathsea/CONFIGURE’.

2.2 Compile-time options

In addition to the configure options listed in the previous section, there are a few things that can be affected at compile-time with C definitions, rather than with configure. Using any of these is unusual.

To specify extra compiler flags (‘-Dname’ in this case), the simplest thing to do is:

     make XCFLAGS="ccoptions"

You can also set the CFLAGS environment variable before running configure. See section “configure environment” in Kpathsea.

Anyway, here are the possibilities:

‘-DFIXPT’
‘-DNO_MF_ASM’

Use the original WEB fixed-point routines for Metafont and MetaPost arithmetic calculations regarding fractions. By default, assembly-language routines are used on x86 hardware with GNU C (unless ‘NO_MF_ASM’ is defined), and floating-point routines are used otherwise.

‘-DIPC_DEBUG’

Report on various interprocess communication activities. See Section 4.6 [IPC and TEX], page 21.

2.3 Additional targets

Web2c has several Make targets besides the standard ones. You can invoke these either in the top level directory of the source distribution (the one containing ‘kpathsea/’ and ‘web2c/’), or in the ‘web2c/’ directory.

‘c-sources’

Make only the C files, translated from the Web sources, presumably because you want to take them to a non-Unix machine.

‘formats’
‘install-formats’

Make or install all the memory dumps (see Section 3.5.2 [Memory dumps], page 10). By default, the standard plain formats plus ‘latex.fmt’ are made. You can add other formats by redefining the fmts, bases, and mems variables. See the top of ‘web2c/Makefile’ for the possibilities.

‘fmts’
‘install-fmts’

Make or install the TEX ‘.fmt’ files. See Section 4.2 [initex invocation], page 16.

‘bases’
‘install-bases’

Make or install the Metafont ‘.base’ files. See Section 5.2 [inimf invocation], page 23.

‘mems’
‘install-mems’

Make or install the MetaPost ‘.mem’ files. See Section 6.2 [inimpost invocation], page 31.

‘triptrap’
‘trip’
‘trap’

‘mptrap’ To run the torture tests for TEX, Metafont, and MetaPost (respectively). See the next section.

2.4 Trip, trap, and mptrap: Torture tests

To validate your TEX, Metafont, and MetaPost executables, run ‘make triptrap’. This runs the trip, trap, and mptrap “torture tests”. See the files ‘triptrap/tripman.tex’,

‘triptrap/trapman.tex’, and ‘triptrap/mptrap.readme’ for detailed information and

background on the tests.

The differences between your executables’ behavior and the standard values will show up on your terminal. The usual differences (these are all acceptable) are:

• string usage and table sizes;

• glue set ratios;

• ‘down4’, ‘right4’, and ‘y4’ commands in DVItype output;

• dates and times.

Any other differences are trouble. The most common culprit in the past has been compiler bugs, especially when optimizing. See section “TEX or Metafont failing” in Kpathsea.

   The files ‘trip.diffs’, ‘mftrap.diffs’, and ‘mptrap.diffs’ in the ‘triptrap’ directory

show the standard diffs against the original output. If you diff your diffs against these files, you should come up clean. For example

     make trip >&mytrip.diffs
     diff triptrap/trip.diffs mytrip.diffs

To run the tests separately, use the targets trip, trap, and mptrap.

To run simple tests for all the programs as well as the torture tests, run ‘make check’. You can compare the output to the distributed file ‘tests/check.log’ if you like.

2.5 Runtime options

Besides the configure- and compile-time options described in the previous sections, you can control a number of parameters (in particular, array sizes) in the ‘texmf.cnf’ runtime file read by Kpathsea (see section “Config files” in Kpathsea).

Rather than exhaustively listing them here, please see the last section of the distributed ‘kpathsea/texmf.cnf’. Some of the more interesting values:

‘main_memory’

Total words of memory available, for TEX, Metafont, and MetaPost. Must remake the format file after changing.

‘extra_mem_bot’

Extra space for “large” TEX data structures: boxes, glue, breakpoints, et al. If you use PiCTEX, you may well want to set this.

‘font_mem_size’

Words of font info available for TEX; this is approximately the total size of all TFM files read.

‘hash_extra’

Additional space for the hash table of control sequence names. Approximately 10,000 control sequences can be stored in the main hash table; if you have a large book with numerous cross-references, this might not be enough, and thus you will want to set hash_extra.

Of course, ideally all arrays would be dynamically expanded as necessary, so the only limiting factor would be the amount of swap space available. Unfortunately, implementing this is extremely difficult, as the fixed size of arrays is assumed in many places throughout the source code. These runtime limits are a practical compromise between the compile-time limits in previous versions, and truly dynamic arrays. (On the other hand, the Web2c BibTEX implementation does do dynamic reallocation of some arrays.)

3 Commonalities

Many aspects of the TEX system are the same among more than one program, so we describe all those pieces together, here.

3.1 Option conventions

To provide a clean and consistent behavior, we chose to have all these programs use the GNU function getopt_long_only to parse command lines.

As a result, you can:

• give the options in any order, interspersed as you wish with non-option arguments;

• use ‘-’ or ‘--’ to start an option name;

• use any unambiguous abbreviation for an option name;

• separate option names and values with either ‘=’ or one or more spaces;

• use filenames that would otherwise look like options by putting them after an option ‘--’.

By convention, non-option arguments, if specified, generally define the name of an input file, as documented for each program.

If a particular option with a value is given more than once, it is the last value that counts.

For example, the following command line specifies the options ‘foo’, ‘bar’, and ‘verbose’; gives the value ‘baz’ to the ‘abc’ option, and the value ‘xyz’ to the ‘quux’ option; and specifies the filename ‘-myfile-’.

     -foo --bar -verb -abc=baz -quux karl --quux xyz -- -myfile-

3.2 Common options

All of these programs accept the standard GNU ‘--help’ and ‘--version’ options, and several programs accept ‘--verbose’. Rather than writing identical descriptions in every node, they are described here.

‘--help’ Print a usage message listing basic usage and all available options to standard output, then exit successfully.

‘--verbose’

Print progress reports to standard output.

‘--version’

Print the version number to standard output, then exit successfully.

TEX, Metafont, and MetaPost have additional options in common:

‘-kpathsea-debug=number’

Set path searching debugging flags according to the bits of number (see section “Debugging” in Kpathsea). You can also specify this in KPATHSEA_DEBUG environment variable (for all Web2c programs). (The command line value overrides.) The most useful value is ‘-1’, to get all available output.

‘-ini’ Enable the “initial” form of the program (see Section 3.5.1 [Initial and virgin], page 9). This is implicitly set if the program name is initex resp. inimf resp.

inimpost.
‘-interaction=string’

Set the interaction mode from the command line. The string must be one of

‘batchmode’, ‘nonstopmode’, ‘scrollmode’, or ‘errorstopmode’.
‘-fmt=dumpname’
‘-base=dumpname’
‘-mem=dumpname’

Use dumpname instead of the program name or a ‘%&’ line to determine the name of the memory dump file read (‘fmt’ for TEX, ‘base’ for Metafont, ‘mem’ for MetaPost). See Section 3.5.2 [Memory dumps], page 10. Also sets the program name to dumpname if no ‘-progname’ option was given.

‘-jobname=string’

Set the job name to string, instead of deriving it from the name of the input file.

‘-parse-first-line’

Check whether the first line of the main input file starts with ‘%&’, and parse it if it does. This line can be used specify the format and/or a TCX file.

‘-progname=string’

Set program (and memory dump) name to string. This may affect the search paths and other values used (see section “Config files” in Kpathsea). Using this option is equivalent to making a link named string to the binary and then invoking the binary under that name. See Section 3.5.2 [Memory dumps], page 10.

‘-recorder’

Enable the filename recorder. This makes the program save a list of the files opened into a file with extension ‘.fls’. For Omega, this option is always on, and the file has extension ‘.ofl’.

‘-translate-file=tcxfile’

Use tcxfile to define which characters are printable and translations between the internal and external character sets. Moreover, tcxfile can be explicitly declared in the first line of the main input file ‘%& -translate-file=tcxfile’. This is the recommended method for portability reasons. See Section 4.5.2 [TCX files], page 19.

‘-file-line-error-style’

Change the way error messages are printed. The alternate style looks like error messages from many compilers and is easier to parse for some editors that drive TEX compilers.

‘-oem’ This option is specific to win32. When used, TEX engines will use the OEM code page rather than the ANSI one to display their messages.

3.3 Path searching

All of the Web2c programs, including TEX, which do path searching use the Kpathsea routines to do so. The precise names of the environment and configuration file variables which get searched for particular file formatted are therefore documented in the Kpathsea manual (see section “Supported file formats” in Kpathsea). Reading ‘texmf.cnf’ (see section “Config files” in Kpathsea), invoking mktex... scripts (see section “mktex scripts” in Kpathsea), and so on are all handled by Kpathsea.

The programs which read fonts make use of another Kpathsea feature: ‘texfonts.map’, which allows arbitrary aliases for the actual names of font files; for example, ‘Times-Roman’ for ‘ptmr8r.tfm’. The distributed (and installed by default) ‘texfonts.map’ includes aliases for many widely available PostScript fonts by their PostScript names.

3.4 Output file location

All the programs generally follow the usual convention for output files. Namely, they are placed in the directory current when the program is run, regardless of any input file location; or, in a few cases, output is to standard output.

For example, if you run ‘tex /tmp/foo’, for example, the output will be in ‘./foo.dvi’

and ‘./foo.log’, not ‘/tmp/foo.dvi’ and ‘/tmp/foo.log’.

However, if the current directory is not writable, the main programs (TEX, Metafont, MetaPost, and BibTEX) make an exception: if the environment variable or config file value TEXMFOUTPUT is set (it is not by default), output files are written to the directory specified. This is useful when you are in some read-only distribution directory, perhaps on a CD-ROM, and want to TEX some documentation, for example.

3.5 Three programs: Metafont, MetaPost, and TEX

TEX, Metafont, and MetaPost have a number of features in common. Besides the ones here, the common command-line options are described in the previous section. The configuration file options that let you control some array sizes and other features are described in Section 2.5 [Runtime options], page 5.

3.5.1 Initial and virgin

The TEX, Metafont, and MetaPost programs each have two main variants, called initial and virgin. As of Web2c 7, one executable suffices for both variants.

The initial form is enabled if:

1. the ‘-ini’ option was specified; or

2. the program name is ‘initex’ resp. ‘inimf’ resp. ‘inimpost’; or

3. the first line of the main input file is ‘%&ini’;

otherwise, the virgin form is used.

The virgin form is the one generally invoked for production use. The first thing it does is read a memory dump (see Section 3.5.2.2 [Determining the memory dump to use], page 11), and then proceeds on with the main job.

The initial form is generally used only to create memory dumps (see the next section). It starts up more slowly than the virgin form, because it must do lengthy initializations that are encapsulated in the memory dump file.

In the past, there was a third form, preloaded executables. This is no longer recommended or widely used; but see the section below if you’re interested anyway. In this case, the memory dump file was read in to the virgin form, a core dump of the running executable was done, and the undump program run to create a new binary. Nowadays, reading memory dumps is fast enough that this is generally no longer worth the cost in disk space and unshared executables.

3.5.1.1 Preloaded executables

Specifying ‘--enable-auto-core’ to configure tells TEX, Metafont, and MetaPost to

suicide with a SIGQUIT on an input filename of ‘HackyInputFileNameForCoreDump.tex’

(all three programs use the ‘.tex’ suffix). This produces a memory dump of the running executable in a file ‘core’. (This is unrelated to the standard memory dump feature in these programs; see Section 3.5.2 [Memory dumps], page 10).

You don’t actually need to do this to produce a core dump. Just typing your quit character (usually ^{⟨}\underline{\overline{CTRL-\}}^{⟩}) when the program is waiting for input (at ‘**’) will have the same result. But a few sites want to reliably generate a core dump without human intervention;

that’s what --enable-auto-core is for.

With the program undump, you can use ‘core’ to reconstitute a preloaded executable, which does not need to read a ‘.fmt’ file to get started. Although preloaded executables save startup time, they have a big disadvantage: neither the disk space to store them nor their code segments (at runtime) can be shared. Therefore, if both tex and latex are running, twice as much memory will be consumed, to the general detriment of performance.

The undump program is not part of the Web2c distribution, but you can get it from the CTAN archives as ‘CTAN:/support/undump’, and it is included in several TEX distributions (see section “unixtex.ftp” in Kpathsea).

3.5.2 Memory dumps

In typical use, TEX, Metafont, and MetaPost require a large number of macros to be predefined; therefore, they support memory dump files, which can be read much more efficiently than ordinary source code.

3.5.2.1 Creating memory dumps

The programs all create memory dumps in slightly idiosyncratic (thought substantially similar) way, so we describe the details in separate sections (references below). The basic idea is to run the initial version of the program (see Section 3.5.1 [Initial and virgin], page 9), read the source file to define the macros, and then execute the \dump primitive.

Also, each program uses a different filename extension for its memory dumps, since although they are completely analogous they are not interchangeable (TEX cannot read a Metafont memory dump, for example).

Here is a list of filename extensions with references to examples of creating memory dumps:

TEX (‘.fmt’) See Section 4.2 [initex invocation], page 16.

Metafont (‘.base’) See Section 5.2 [inimf invocation], page 23.

MetaPost (‘.mem’) See Section 6.2 [inimpost invocation], page 31.

When making memory dumps, the programs read environment variables and configuration files for path searching and other values as usual. If you are making a new installation and have environment variables pointing to an old one, for example, you will probably run into difficulties.

3.5.2.2 Determining the memory dump to use

The virgin form (see Section 3.5.1 [Initial and virgin], page 9) of each program always reads a memory dump before processing normal source input. All three programs determine the memory dump to use in the same way:

1. If the first non-option command-line argument begins with ‘&’, the program uses the remainder of that argument as the memory dump name. For example, running ‘tex \&super’ reads ‘super.fmt’. (The backslash protects the ‘&’ against interpretation by the shell.)

2. If the ‘-fmt’ resp. ‘-base’ resp. ‘-mem’ option is specified, its value is used.

3. If the ‘-progname’ option is specified, its value is used.

4. If the first line of the main input file (which must be specified on the command line, not in response to ‘**’) is %&dump, and dump is an existing memory dump of the appropriate type, dump is used. The first line of the main input file can also specify which character translation file is to be used: %&-translate-file=tcxfile (see Section 4.5.2 [TCX files], page 19). These two roles can be combined: %&dump -translate-file=tcxfile. If this is done, the name of the dump must be given first.

5. Otherwise, the program uses the program invocation name, most commonly ‘tex’ resp. ‘mf’ resp. ‘mpost’. For example, if ‘latex’ is a link to ‘tex’, and the user runs ‘latex foo’, ‘latex.fmt’ will be used.

3.5.2.3 Hardware and memory dumps

By default, memory dump files are generally sharable between architectures of different types; specifically, on machines of different endianness (see section “Byte order” in GNU C Library). (This is a feature of the Web2c implementation, and is not true of all TEX implementations.) If you specify ‘--disable-dump-share’ to configure, however, memory dumps will be endian-dependent.

The reason to do this is speed. To achieve endian-independence, the reading of memory dumps on LittleEndian architectures, such as PC’s and DEC architectures, is somewhat slowed (all the multibyte values have to be swapped). Usually, this is not noticeable, and the advantage of being able to share memory dumps across all platforms at a site far outweighs the speed loss. But if you’re installing Web2c for use on LittleEndian machines only, perhaps on a PC being used only by you, you may wish to get maximum speed.

TEXnically, even without ‘--disable-dump-share’, sharing of ‘.fmt’ files cannot be guaranteed to work. Floating-point values are always written in native format, and hence will generally not be readable across platforms. Fortunately, TEX uses floating point only to represent glue ratios, and all common formats (plain, LaTEX, AMSTEX, …) do not do any glue setting at ‘.fmt’-creation time. Metafont and MetaPost do not use floating point in any dumped value at all.

Incidentally, different memory dump files will never compare equal byte-for-byte, because the program always dumps the current date and time. So don’t be alarmed by just a few bytes difference.

If you don’t know what endianness your machine is, and you’re curious, here is a little C program to tell you. (The configure script contains a similar program.) This is from the book C: A Reference Manual, by Samuel P. Harbison and Guy L. Steele Jr. (see Appendix B [References], page 55).

     main ()
     {
       /* Are we little or big endian?  From Harbison&Steele.  */
       union
       {
         long l;
         char c[sizeof (long)];
       } u;
       u.l = 1;
       if (u.c[0] == 1)
         printf ("LittleEndian\n");
       else if (u.c[sizeof (long) - 1] == 1)
         printf ("BigEndian\n");
       else
         printf ("unknownEndian");
       exit (u.c[sizeof (long) - 1] == 1);
     }

3.5.3 Editor invocation

TEX, Metafont, and MetaPost all (by default) stop and ask for user intervention at an error. If the user responds with e or E, the program invokes an editor.

Specifying ‘--with-editor=cmd’ to configure sets the default editor command string to cmd. The environment variables/configuration values TEXEDIT, MFEDIT, and MPEDIT (respectively) override this. If ‘--with-editor’ is not specified, the default is vi +%d %s.

In this string, ‘%d’ is replaced by the line number of the error, and ‘%s’ is replaced by the name of the current input file.

3.5.4 \input filenames

TEX, Metafont, and MetaPost source programs can all read other source files with the \input (TEX) and input (MF and MP) primitives:

     \input name % in TeX

The file name can always be terminated with whitespace; for Metafont and MetaPost, the statement terminator ‘;’ also works. (LaTEX and other macro packages provide other interfaces to \input that allow different notation; here we are concerned only with the primitive operation.) This means that \input filenames cannot directly contain whitespace, even though Unix has no trouble. Sorry.

On the other hand, various C library routines and Unix itself use the null byte (character code zero, ASCII NUL) to terminate strings. So filenames in Web2c cannot contain nulls, even though TEX itself does not treat NUL specially.

Furthermore, some older Unix variants do not allow eight-bit characters (codes 128–255) in filenames.

For maximal portability of your document across systems, use only the characters ‘a’– ‘z’, ‘0’–‘9’, and ‘.’, and restrict your filenames to at most eight characters (not including the extension), and at most a three-character extension. Do not use anything but simple filenames, since directory separators vary among systems; instead, add the necessary directories to the appropriate search path.

Finally, the present Web2c implementation does ‘~’ and ‘$’ expansion on name, unlike Knuth’s original implementation and older versions of Web2c. Thus:

     \input ~jsmith/$foo.bar

will dereference the environment variable or Kpathsea config file value ‘foo’ and read that file extended with ‘.bar’ in user ‘jsmith’’s home directory. (You can also use braces, as in ‘${foo}bar’ if you want to follow the variable name with a letter, numeral, or ‘_’.)

(So you could define an environment variable value including whitespace and get the program to read such a filename that way, if you need to.)

In all the common TEX formats (plain TEX, LaTEX, AMSTEX), the characters ‘~’ and ‘~’ have special category codes, so to actually use these in a document you have to change their catcodes or use \string. (The result is unportable anyway, see the suggestions above.) The place where they are most likely to be useful is when typing interactively.

4 TEX: Typesetting

TEX is a typesetting system: it was especially designed to handle complex mathematics, as well as most ordinary text typesetting.

TEX is a batch language, like C or Pascal, and not an interactive “word processor”: you compile a TEX input file into a corresponding device-independent (DVI) file (and then translate the DVI file to the commands for a particular output device). This approach has both considerable disadvantages and considerable advantages. For a complete description of the TEX language, see The TEXbook (see Appendix B [References], page 55). Many other books on TEX, introductory and otherwise, are available.

4.1 tex invocation

TEX (usually invoked as tex) formats the given text and commands, and outputs a corresponding device-independent representation of the typeset document. This section merely describes the options available in the Web2c implementation. For a complete description of the TEX typesetting language, see The TEXbook (see Appendix B [References], page 55).

TEX, Metafont, and MetaPost process the command line (described here) and determine their memory dump (fmt) file in the same way (see Section 3.5.2 [Memory dumps], page 10). Synopses:

     tex [option]... [texname[.tex]] [tex-commands]
     tex [option]... \first-line
     tex [option]... &fmt args

TEX searches the usual places for the main input file texname (see section “Supported file formats” in Kpathsea), extending texname with ‘.tex’ if necessary. To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

After texname is read, TEX processes any remaining tex-commands on the command line as regular TEX input. Also, if the first non-option argument begins with a TEX escape character (usually \), TEX processes all non-option command-line arguments as a line of regular TEX input.

If no arguments or options are specified, TEX prompts for an input file name with ‘**’. TEX writes the main DVI output to the file ‘basetexname.dvi’, where basetexname is

the basename of texname, or ‘texput’ if no input file was specified. A DVI file is a deviceindependent binary representation of your TEX document. The idea is that after running TEX, you translate the DVI file using a separate program to the commands for a particular output device, such as a PostScript printer (see section “Introduction” in Dvips) or an X Window System display (see xdvi(1)).

TEX also reads TFM files for any fonts you load in your document with the \font primitive. By default, it runs an external program named ‘mktextfm’ to create any nonexistent TFM files. You can disable this at configure-time or runtime (see section “mktex configuration” in Kpathsea). This is enabled mostly for the sake of the EC fonts, which can be generated at any size.

TEX can write output files, via the \openout primitive; this opens a security hole vulnerable to Trojan horse attack: an unwitting user could run a TEX program that overwrites, say, ‘~/.rhosts’. (MetaPost has a write primitive with similar implications). To alleviate this, there is a configuration variable openout_any, which selects one of three levels of security. When it is set to ‘a’ (for “any”), no restrictions are imposed. When it is set to ‘r’ (for “restricted”), filenames beginning with ‘.’ are disallowed (except ‘.tex’ because LaTEX needs it). When it is set to ‘p’ (for “paranoid”) additional restrictions are imposed: an absolute filename must refer to a file in (a subdirectory) of TEXMFOUTPUT, and any attempt to go up a directory level is forbidden (that is, paths may not contain a ‘..’ component). The paranoid setting is the default. (For backwards compatibility, ‘y’ and ‘1’ are synonyms of ‘a’, while ‘n’ and ‘0’ are synonyms for ‘r’.)

In any case, all \openout filenames are recorded in the log file, except those opened on the first line of input, which is processed when the log file has not yet been opened. (If you as a TEX administrator wish to implement more stringent rules on \openout, modifying the function openoutnameok in ‘web2c/lib/texmfmp.c’ is intended to suffice.)

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-kpathsea-debug=number’
‘-ini’
‘-fmt=fmtname’
‘-progname=string’
‘-translate-file=tcxfile’

These options are common to TEX, Metafont, and MetaPost. See Section 3.2 [Common options], page 7.

‘-ipc’
‘-ipc-start’

With either option, TEX writes its DVI output to a socket as well as to the usual ‘.dvi’ file. With ‘-ipc-start’, TEX also opens a server program at the other end to read the output. See Section 4.6 [IPC and TEX], page 21. These options are available only if the ‘--enable-ipc’ option was specified to configure during installation of Web2c.

‘-mktex=filetype’
‘-no-mktex=filetype’

Turn on or off the ‘mktex’ script associated with filetype. The only values that make sense for filetype are ‘tex’ and ‘tfm’,

‘-mltex’ If INITEX (see Section 3.5.1 [Initial and virgin], page 9), enable MLTEX extensions such as \charsubdef. Implicitly set if the program name is mltex. See Section 4.5.1 [MLTEX], page 18.

‘-output-comment=string’

Use string as the DVI file comment. Ordinarily, this comment records the date and time of the TEX run, but if you are doing regression testing, you may not want the DVI file to have this spurious difference. This is also taken from the environment variable and config file value ‘output_comment’.

‘-shell-escape’

Enable the ‘\write18{shell-command}’ feature. This is also enabled if the environment variable or config file value ‘shell_escape’ is set to ‘t’. (For backwards compatibility, ‘y’ and ‘1’ are accepted as synonyms of ‘t’). It is disabled by default to avoid security problems. When enabled, the shell-command string (which first undergoes the usual TEX expansions, just as in ‘\special’) is passed to the command shell (via the C library function ‘system’). The output of shell-command is not diverted anywhere, so it will not appear in the log file. The system call either happens at ‘\output’ time or right away, according to the absence or presence of the ‘\immediate’ prefix, as usual for \write. (If you as a TEX administrator wish to implement more stringent rules on what can be executed, you will need to modify ‘tex.ch’.)

‘-src-specials’
‘-src-specials=string’

This option makes TEX output specific source information using ‘\special’ commands in the DVI file. These ‘\special’ track the current file name and line number. Using the first form of this option, the ‘\special’ are inserted automatically. In the second form of the option, string is a comma separated list of the following values: ‘cr’, ‘display’, ‘hbox’, ‘math’, ‘par’, ‘parend’, ‘vbox’. You can use this list to specify where you want TEX to output such commands. By example, ‘-src-specials=cr,math’ will output source information every line and every math formula. These commands can be used with the appropriate DVI viewer and text editor to switch from the current position in the editor to the same position in the viewer and back from the viewer to the editor. Note that this option works by inserting ‘\special’ commands into the token stream, and that these additional tokens can in principle be recovered by the right macros. If you run across a case, let us know, because this counts as a bug. However, these bugs are very hard to fix without making significant changes to TEX, so please don’t count on them being fixed. Redefining ‘\special’ will not affect the functioning of this option. The commands we insert into the token stream have been hard-coded to always be the ‘\special’ primitive. Also note that TEX does not pass the trip test when this option is enabled.

4.2 initex invocation

initex is the “initial” form of TEX, which does lengthy initializations avoided by the “virgin” (vir) form, so as to be capable of dumping ‘.fmt’ files (see Section 3.5.2 [Memory dumps], page 10). For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 4.1 [tex invocation], page 14. Unlike Metafont and MetaPost, many format files are commonly used with TEX. The

standard one implementing the features described in the TEXbook is ‘plain.fmt’, also known as ‘tex.fmt’ (again, see Section 3.5.2 [Memory dumps], page 10). It is created by default during installation, but you can also do so by hand if necessary (e.g., if an update to ‘plain.tex’ is issued):

     initex '\input plain \dump'

(The quotes prevent interpretation of the backslashes from the shell.) Then install the

resulting ‘plain.fmt’ in ‘$(fmtdir)’ (‘/usr/local/share/texmf/web2c’ by default), and

link ‘tex.fmt’ to it.

The necessary invocation for generating a format file differs for each format, so instructions that come with the format should explain. The top-level ‘web2c’ Makefile has targets for making most common formats: plain latex amstex texinfo eplain. See Section 4.4 [Formats], page 17, for more details on TEX formats.

4.3 virtex invocation

virtex is the “virgin” form of TEX, which avoids the lengthy initializations done by the “initial” (ini) form, and is thus what is generally used for production work. For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 4.1 [tex invocation], page 14.

4.4 Formats

TEX formats are large collections of macros, possibly dumped into a ‘.fmt’ file (see Section 3.5.2 [Memory dumps], page 10) by initex (see Section 4.2 [initex invocation], page 16). A number of formats are in reasonably widespread use, and the Web2c Makefile has targets to make the versions current at the time of release. You can change which formats are automatically built by setting the fmts Make variable; by default, only the ‘plain’ and ‘latex’ formats are made.

You can get the latest versions of most of these formats from the CTAN archives in subdirectories of ‘CTAN:/macros’ (for CTAN info, see section “unixtex.ftp” in Kpathsea). The archive ftp://ftp.tug.org/tex/lib.tar.gz (also available from CTAN) contains most of these formats (although perhaps not the absolute latest version), among other things.

latex The most widely used format. The current release is named ‘LaTEX 2e’; new versions are released approximately every six months, with patches issued as needed. The old release was called ‘LaTEX 2.09’, and is no longer maintained or supported. LaTEX attempts to provide generic markup instructions, such as “emphasize”, instead of specific typesetting instructions, such as “use the 10 pt Computer Modern italic font”.

amstex The official typesetting system of the American Mathematical Society, used to produce nearly all of its publications, e.g., Mathematical Reviews. Like LaTEX, it encourages generic markup commands. The AMS also provides a LaTEX package for authors who prefer LaTEX (see the ‘amslatex’ item below).

texinfo The documentation system developed and maintained by the Free Software Foundation for their software manuals. It can be automatically converted into plain text, a machine-readable on-line format called ‘info’, HTML, etc. eplain The “expanded plain” format provides various common features (e.g., sym-

bolic cross-referencing, tables of contents, indexing, citations using BibTEX), for those authors who prefer to handle their own high-level formatting.

lamstex Augments AMSTEX with LaTEX-like features.

amslatex An LaTEX package (see ‘latex’ item above), that augments LaTEX with AMSTEX-like features.

slitex An obsolete LaTEX 2.09 format for making slides. It is replaced by the ‘slides’ document class.

4.5 Languages and hyphenation

TEX supports most natural languages. See also Section 4.7 [TEX extensions], page 21.

4.5.1 MLTEX: Multi-lingual TEX

Multi-lingual TEX (mltex) is an extension of TEX originally written by Michael Ferguson and now updated and maintained by Bernd Raichle. It allows the use of non-existing glyphs in a font by declaring glyph substitutions. These are restricted to substitutions of an accented character glyph, which need not be defined in the current font, by its appropriate \accent construction using a base and accent character glyph, which do have to exist in the current font. This substitution is automatically done behind the scenes, if necessary, and thus MLTEX additionally supports hyphenation of words containing an accented character glyph for fonts missing this glyph (e.g., Computer Modern). Standard TEX suppresses hyphenation in this case.

MLTEX works at ‘.fmt’-creation time: the basic idea is to specify the ‘-mltex’ option to TEX when you \dump a format. Then, when you subsequently invoke TEX and read that .fmt file, the MLTEX features described below will be enabled.

Generally, you use special macro files to create an MLTEX .fmt file. See:

     CTAN:/systems/generic/mltex
     ftp://ftp.univ-rennes1.fr/pub/GUTenberg/french/

The sections below describe the two new primitives that MLTEX defines. Aside from these, MLTEX is completely compatible with standard TEX.

4.5.1.1 \charsubdef: Character substitutions

The most important primitive MLTEX adds is \charsubdef, used in a way reminiscent

of \chardef:
     \charsubdef composite [=] accent base

Each of composite, accent, and base are font glyph numbers, expressed in the usual TEX syntax: `\e symbolically, '145 for octal, "65 for hex, 101 for decimal.

MLTEX’s \charsubdef declares how to construct an accented character glyph (not necessarily existing in the current font) using two character glyphs (that do exist). Thus it defines whether a character glyph code, either typed as a single character or using the \char primitive, will be mapped to a font glyph or to an \accent glyph construction.

For example, if you assume glyph code 138 (decimal) for an e-circumflex (ê) and you are using the Computer Modern fonts, which have the circumflex accent in position 18 and lowercase ‘e’ in the usual ASCII position 101 decimal, you would use \charsubdef as follows:

     \charsubdef 138 = 18 101

For the plain TEX format to make use of this substitution, you have to redefine the circumflex accent macro \^in such a way that if its argument is character ‘e’ the expansion \char138 is used instead of \accent18 e. Similar \charsubdef declaration and macro redefinitions have to be done for all other accented characters.

To disable a previous \charsubdef c, redefine c as a pair of zeros. For example:

     \charsubdef '321 = 0 0  % disable N tilde

(Octal '321 is the ISO Latin-1 value for the Spanish N tilde.)

\charsubdef commands should only be given once. Although in principle you can use \charsubdef at any time, the result is unspecified. If \charsubdef declarations are changed, usually either incorrect character dimensions will be used or MLTEX will output missing character warnings. (The substitution of a \charsubdef is used by TEX when appending the character node to the current horizontal list, to compute the width of a horizontal box when the box gets packed, and when building the \accent construction at \shipout-time. In summary, the substitution is accessed often, so changing it is not desirable, nor generally useful.)

4.5.1.2 \tracingcharsubdef: Substitution diagnostics

To help diagnose problems with ‘\charsubdef’, MLTEX provides a new primitive parameter, \tracingcharsubdef. If positive, every use of \charsubdef will be reported. This can help track down when a character is redefined.

In addition, if the TEX parameter \tracinglostchars is 100 or more, the character substitutions actually performed at \shipout-time will be recorded.

4.5.2 TCX files: Character translations

TCX (TEX character translation) files help TEX support direct input of 8-bit international characters if fonts containing those characters are being used. Specifically, they map an input (keyboard) character code to the internal TEX character code (a superset of ASCII).

Of the various proposals for handling more than one input encoding, TCX files were chosen because they follow Knuth’s original ideas for the use of the ‘xhcr’ and ‘xord’ tables. He ventured that these would be changed in the WEB source in order to adjust the actual version to a given environment. It turned out, however, that recompiling the WEB sources is not as simple task as Knuth predicted; therefore, TCX files, providing the possibility of changing of the conversion tables on on-the-fly, has been implemented instead.

This approach limits the portability of TEX documents, as some implementations do not support it (or use a different method for input-internal reencoding). It may also be problematic to determine the encoding to use for a TEX document of unknown provenance; in the worst case, failure to do so correctly may result in subtle errors in the typeset output.

While TCX files can be used with any format, using them breaks the LaTEX ‘inputenc’ package. This is why you should either use tcxfile or ‘inputenc’ in LaTEX files, but never both.

This is entirely independent of the MLTEX extension (see Section 4.5.1 [MLTeX], page 18): whereas a TCX file defines how an input keyboard character is mapped to TEX’s internal code, MLTEX defines substitutions for a non-existing character glyph in a font with a \accent construction made out of two separate character glyphs. TCX files involve no new primitives; it is not possible to specify that an input (keyboard) character maps to more than one character.

Specifying TCX files:

• You can specify a TCX file to be used for a particular TEX run by specifying the command-line option ‘-translate-file=tcxfile’ or (preferably) specifying it explicitly in the first line of the main document ‘%& -translate-file=tcxfile’.

• TCX files are searched for along the WEB2C path.

• INITEX ignores TCX files.

The Web2c distribution comes with at least two TCX files, ‘il1-t1.tcx’ and ‘il2-t1.tcx’. These support ISO Latin 1 and ISO Latin 2, respectively, with Cork-encoded fonts (a.k.a. the T1 encoding). TCX files for Czech, Polish, and Slovak are also provided.

Syntax of TCX files:

1. Line-oriented. Blank lines are ignored.

2. Whitespace is ignored except as a separator.

3. Comments start with ‘%’ and continue to the end of the line.

4. Otherwise, a line consists of one or two character codes:

         src [dest]

5. Each character code may be specified in octal with a leading ‘0’, hexadecimal with a leading ‘0x’, or decimal otherwise. Values must be between 0 and 255, inclusive (decimal).

6. If the dest code is not specified, it is taken to be the same as src.

7. If the same src code is specified more than once, it is the last definition that counts.

Finally, here’s what happens: when TEX sees an input character with code src, it 1) changes src to dest; and 2) makes code the dest “printable”, i.e., printed as-is in diagnostics and the log file instead of in ‘^^’ notation.

By default, no characters are translated, and character codes between 32 and 126 inclusive (decimal) are printable. It is not possible to make these (or any) characters unprintable.

Specifying translations for the printable ASCII characters (codes 32–127) will yield unpredictable results. Additionally you shouldn’t make the following characters printable: ^^I (TAB), ^^J (line feed), ^^M (carriage return), and ^^? (delete), since TEX uses them in various ways.

Thus, the idea is to specify the input (keyboard) character code for src, and the output (font) character code for dest.

4.5.3 Patgen: Creating hyphenation patterns

Patgen creates hyphenation patterns from dictionary files for use with TEX. Synopsis:

     patgen dictionary patterns output translate

Each argument is a filename. No path searching is done. The output is written to the file output.

In addition, Patgen prompts interactively for other values.

For more information, see Word hy-phen-a-tion by com-puter by Frank Liang (see Appendix B [References], page 55), and also the ‘patgen.web’ source file.

The only options are ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7).

4.6 IPC and TEX

(Sorry, but I’m not going to write this unless someone actually uses this feature. Let me know.)

This functionality is available only if the ‘--enable-ipc’ option was specified to configure during installation of Web2c (see Chapter 2 [Installation], page 2).

If you define IPC_DEBUG before compilation (e.g., with ‘make XCFLAGS=-DIPC_DEBUG’), TEX will print messages to standard error about its socket operations. This may be helpful if you are, well, debugging.

4.7 TEX extensions

The base TEX program has been extended in many ways. Here’s a partial list. Please send information on extensions not listed here to the address in section “Reporting bugs” in Kpathsea.

e-TEX Adds many new primitives, including right-to-left typesetting. Available from

          http://www.vms.rhbnc.ac.uk/e-TeX/ and ‘CTAN:/systems/e-tex’.

Omega Adds Unicode support, right-to-left typesetting, and more. Available from

          http://www.ens.fr/omega and ‘CTAN:/systems/omega’.

pdfTEX A variant of TEX that produces PDF instead of DVI files. It also includes primitives for hypertext. Available from ‘CTAN:/systems/pdftex’.

‘TeX--XeT’

Adds primitives and DVI opcodes for right-to-left typesetting (as used in Arabic, for example). An old version for TEX 3.1415 is available from ‘CTAN:/systems/knuth/tex--xet’. A newer version is included in e-TEX.

File-handling TEX Adds primitives for creating multiple DVI files in a single run; and appending to output files as well as overwriting. Web2c implementation available in the

          distribution file ‘web2c/contrib/file-handling-tex’.

5 Metafont: Creating typeface families

Metafont is a system for producing shapes; it was designed for producing complete typeface families, but it can also produce geometric designs, dingbats, etc. And it has considerable mathematical and equation-solving capabilities which can be useful entirely on their own.

Metafont is a batch language, like C or Pascal: you compile a Metafont program into a corresponding font, rather than interactively drawing lines or curves. This approach has both considerable disadvantages (people unfamiliar with conventional programming languages will be unlikely to find it usable) and considerable advantages (you can make your design intentions specific and parameterizable). For a complete description of the Metafont language, see The METAFONTbook (see Appendix B [References], page 55).

5.1 mf invocation

Metafont (usually invoked as mf) reads character definitions specified in the Metafont programming language, and outputs the corresponding font. This section merely describes the options available in the Web2c implementation. For a complete description of the Metafont language, see The Metafontbook (see Appendix B [References], page 55).

Metafont processes its command line and determines its memory dump (base) file in a way exactly analogous to MetaPost and TEX (see Section 4.1 [tex invocation], page 14, and see Section 3.5.2 [Memory dumps], page 10). Synopses:

     mf [option]... [mfname[.mf]] [mf-commands]
     mf [option]... \first-line
     mf [option]... &base args

Most commonly, a Metafont invocation looks like this:

     mf '\mode:=mode; mag:=magnification; input mfname'

(The single quotes avoid unwanted interpretation by the shell.)

Metafont searches the usual places for the main input file mfname (see section “Supported file formats” in Kpathsea), extending mfname with ‘.mf’ if necessary. To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program. By default, Metafont runs an external program named ‘mktexmf’ to create any nonexistent Metafont source files you input. You can disable this at configure-time or runtime (see section “mktex configuration” in Kpathsea). This is mostly for the sake of the EC fonts, which can be generated at any size.

Metafont writes the main GF output to the file ‘basemfname.nnngf’, where nnn is the font resolution in pixels per inch, and basemfname is the basename of mfname, or ‘mfput’ if no input file was specified. A GF file contains bitmaps of the actual character shapes. Usually GF files are converted immediately to PK files with GFtoPK (see Section 10.2 [gftopk invocation], page 46), since PK files contain equivalent information, but are more compact. (Metafont output in GF format rather than PK for only historical reasons.)

Metafont also usually writes a metric file in TFM format to ‘basemfname.tfm’. A TFM file contains character dimensions, kerns, and ligatures, and spacing parameters. TEX reads only this .tfm file, not the GF file.

The mode in the example command above is a name referring to a device definition (see Section 5.4 [Modes], page 24); for example, localfont or ljfour. These device definitions must generally be precompiled into the base file. If you leave this out, the default is proof mode, as stated in The Metafontbook, in which Metafont outputs at a resolution of 2602 dpi; this is usually not what you want. The remedy is simply to assign a different mode—localfont, for example.

The magnification assignment in the example command above is a magnification factor; for example, if the device is 600 dpi and you specify mag:=2, Metafont will produce output at 1200 dpi. Very often, the magnification is an expression such as magstep(.5), corresponding to a TEX “magstep”, which are factors of 1.2\sqrt{2}.

After running Metafont, you can use the font in a TEX document as usual. For example:

     \font\myfont = newfont
     \myfont Now I am typesetting in my new font (minimum hamburgers).

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-kpathsea-debug=number’
‘-ini’
‘-base=basename’
‘-progname=string’
‘-translate-file=tcxfile’

These options are common to TEX, Metafont, and MetaPost. See Section 3.2 [Common options], page 7.

‘-mktex=filetype’
‘-no-mktex=filetype’

Turn on or off the ‘mktex’ script associated with filetype. The only value that makes sense for filetype is ‘mf’.

5.2 inimf invocation

inimf is the “initial” form of Metafont, which does lengthy initializations avoided by the “virgin” (vir) form, so as to be capable of dumping ‘.base’ files (see Section 3.5.2 [Memory dumps], page 10). For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 5.1 [mf invocation], page 22. The only memory dump file commonly used with Metafont is the default ‘plain.base’,

also known as ‘mf.base’ (again, see Section 3.5.2 [Memory dumps], page 10). It is created by default during installation, but you can also do so by hand if necessary (e.g., if a Metafont update is issued):

     inimf '\input plain; input modes; dump'

(The quotes prevent interpretation of the backslashes from the shell.) Then install the

resulting ‘plain.base’ in ‘$(basedir)’ (‘/usr/local/share/texmf/web2c’ by default),

and link ‘mf.base’ to it.

For an explanation of the additional ‘modes.mf’ file, see Section 5.4 [Modes], page 24. This file has no counterpart in TEX or MetaPost.

In the past, it was sometimes useful to create a base file ‘cmmf.base’ (a.k.a. ‘cm.base’), with the Computer Modern macros also included in the base file. Nowadays, however, the additional time required to read ‘cmbase.mf’ is exceedingly small, usually not enough to be worth the administrative hassle of updating the ‘cmmf.base’ file when you install a new version of ‘modes.mf’. People actually working on a typeface may still find it worthwhile to create their own base file, of course.

5.3 virmf invocation

virmf is the “virgin” form of Metafont, which avoids the lengthy initializations done by the “initial” (ini) form, and is thus what is generally used for production work. Usually it is invoked under the name ‘mf’. For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 5.1 [mf invocation], page 22.

5.4 Modes: Device definitions for Metafont

Running Metafont and creating Metafont base files requires information that TEX and MetaPost do not: mode definitions which specify device characteristics, so Metafont can properly rasterize the shapes.

When making a base file, a file containing modes for locally-available devices should be

input after ‘plain.mf’. One commonly used file is ftp://ftp.tug.org/tex/modes.mf; it

includes all known definitions.

If, however, for some reason you have decreased the memory available in your Metafont, you may need to copy ‘modes.mf’ and remove the definitions irrelevant to you (probably most of them) instead of using it directly. (Or, if you’re a Metafont hacker, maybe you can suggest a way to redefine mode_def and/or mode_setup; right now, the amount of memory used is approximately four times the total length of the mode_def names, and that’s a lot.)

If you have a device not included in ‘modes.mf’, please see comments in that file for how to create the new definition, and please send the definition to tex-fonts@mail.tug.org to get it included in the next release of ‘modes.mf’.

Usually, when you run Metafont you must supply the name of a mode that was dumped in the base file. But you can also define the mode characteristics dynamically, by invoking Metafont with an assignment to smode instead of mode, like this:

     mf '\smode:="newmode.mf"; mag:=magnification; input mfname'

This is most useful when you are working on the definition of a new mode.

The magnification and mfname arguments are explained in Section 5.1 [mf invocation], page 22. In the file ‘newmode.mf’, you should have the following (with no mode_def or enddef), if you are using ‘modes.mf’ conventions:

     mode_param (pixels_per_inch, dpi);
     mode_param (blacker, b);
     mode_param (fillin, f);
     mode_param (o_correction, o);
     mode_common_setup_;

(Of course, you should use real numbers for dpi, b, f, and o.)

For more information on the use of smode, or if you are not using ‘modes.mf’, see page 269 of The Metafontbook.

5.5 Online Metafont graphics

The Web2c implementation of Metafont can do online graphics with a number of devices. (See the Metafont manual for more information about how to draw on your screen.) By default, no graphics support is enabled.

Metafont examines the MFTERM environment variable or config file value at runtime, or the TERM environment variable if MFTERM is not set, to determine the device support to use. Naturally, only the devices for which support has been compiled in can be selected.

Here is a table of the possibilities, showing the MFTERM value and the corresponding configure option(s) in parentheses.

epsf (‘--with-epsfwin’) Encapsulated PostScript pseudo-window server (see ‘web2c/window/epsf.c’). This device produces an EPS file containing the graphics which would be displayed online on other devices. The name of the EPS file defaults to metafont.eps but can be changed by setting the MFEPSF environment variable to the new filename. Contributed by Mathias Herberts.

hp2627 (‘--with-hp2627win’) HP2627a color graphics terminals.

mftalk  (‘--with-mftalkwin’) Generic window server (see ‘web2c/window/mftalk.c’).

next (‘--with-next’) NeXT window system. This requires a separate program, called DrawingServant, available separately. See the ‘web2c/window/next.c’.

regis  (‘--with-regiswin’) Regis terminals.

sun (‘--with-suntoolswin’) The old Suntools (not any flavor of X) window system. (You can get the even older SunWindows gfx system by using ‘sun-gfx.c’.)

tek (‘--with-tektronixwin’) Tektronix terminals.

uniterm (‘--with-unitermwin’) Uniterm, Simon Poole’s emulator of a smart Tektronix 4014 terminal. This may work with regular Tektronix terminals as well; it’s faster than the driver ‘--with-tek’ selects.

xterm  (‘--with-x11win’, ‘--with-x’, ‘--with-x11’) The X window system (version

11).

There are two variants of the X11 support, one that works with the Xt toolkit, and another that works directly with Xlib. The Xt support is more efficient and has more functionality, so it is the default. If you must use the Xlib support,

          use ‘configure --with-x --with-x-toolkit=no’.

You cannot specify any of the usual X options (e.g., ‘-geometry’) on the Metafont command line, but you can specify X resources in your ‘~/.Xdefaults’ or ‘~/.Xresources’ file. The class name is Metafont. If you’re using the Xt support, all the usual X toolkit resources are supported. If you’re using the Xlib support, only the geometry resource is supported. You specify the X display to which Metafont connects in the DISPLAY environment variable, as usual.

Writing support for a new device is straightforward. Aside from defining the basic drawing routines that Metafont uses (see ‘mf.web’), you only have to add another entry to the tables on the last page of ‘web2c/lib/texmfmp.c’. Or you can write an independent program and use MFtalk (see ‘web2c/window/mftalk.c’).

5.6 GFtoDVI: Character proofs of fonts

GFtoDVI makes proof sheets from a GF bitmap file as output by, for example, Metafont (see Chapter 5 [Metafont], page 22). This is an indispensable aid for font designers or Metafont hackers. Synopsis:

     gftodvi [option]... gfname[gf]

The font gfname is searched for in the usual places (see section “Glyph lookup” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

The suffix ‘gf’ is supplied if not already present. This suffix is not an extension; no ‘.’ precedes it: for instance ‘cmr10.600gf’.

The output filename is the basename of gfname extended with ‘.dvi’, e.g., ‘gftodvi

/wherever/foo.600gf’ creates ‘./foo.dvi’.

The characters from gfname appear one per page in the DVI output, with labels, titles, and annotations, as specified in Appendix H (Hardcopy Proofs) of The Metafontbook.

GFtoDVI uses several fonts besides gfname itself:

• gray font (default ‘gray’): for the pixels that actually make up the character. Simply using black is not right, since then labels, key points, and other information could not be shown.

• title font (default ‘cmr8’): for the header information at the top of each output page.

• label font (default ‘cmtt10’): for the labels on key points of the figure.

• slant font (no default): for diagonal lines, which are otherwise simulated using horizontal and vertical rules.

To change the default fonts, you must use special commands in your Metafont source file.

The program accepts the following option, as well as the standard ‘-verbose’, ‘-help’, and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-overflow-label-offset=points’

Typeset the so-called overflow labels, if any, points TEX points from the right edge of the character bounding box. The default is a little over two inches (ten million scaled points, to be precise). Overflow equations are used to locate coordinates when their actual position is too crowded with other information.

5.7 MFT: Prettyprinting Metafont source

MFT translates a Metafont program into a TEX document suitable for typesetting, with the aid of TEX macros defined in the file ‘mftmac.tex’. Synopsis:

 mft [option]... mfname[.mf]

MFT searches the usual places for mfname (see section “Supported file formats” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program. The output goes to the basename of mfname extended

with ‘.tex’, e.g., ‘mft /wherever/foo.mf’ creates ‘./foo.tex’.

Line breaks in the input are carried over into the output; moreover, blank spaces at the beginning of a line are converted to quads of indentation in the output. Thus, you have full control over the indentation and line breaks. Each line of input is translated independently of the others.

Further control is allowed via Metafont comments:

• Metafont comments following a single ‘%’ should be valid TEX input. But Metafont material can be included within vertical bars in a comment; this will be translated by MFT as if it were regular Metafont code. For example, a comment like ‘% |x2r| is the tip of the bowl’ will be translated into the TEX ‘% $x_{2r}$ is the ...’, i.e., the ‘x2r’ is treated as an identifier.

• ‘%%’ indicates that the remainder of an input line should be copied verbatim to the output. This is typically used to introduce additional TEX material at the beginning or an MFT job, e.g. code to modify the standard layout or the formatting macros defined in ‘mftmac.tex’, or to add a line saying ‘%%\bye’ at the end of the job. (MFT doesn’t add this automatically in order to allow processing several files produces by MFT in the same TEX job.)

• ‘%%% token1 other-tokens’ introduces a change in MFT’s formatting rules; all the other-tokens will henceforth be translated according to the current conventions for token1. The tokens must be symbolic (i.e., not numeric or string tokens). For example, the input line

     %%% addto fill draw filldraw

says to format the ‘fill’, ‘draw’, and ‘filldraw’ operations of plain Metafont just like the primitive token ‘addto’, i.e., in boldface type. Without such reformatting commands, MFT would treat ‘fill’ like an ordinary tag or variable name. In fact, you need a ‘%%%’ command even to get parentheses to act like delimiters.

• ‘%%%%’ introduces an MFT comment, i.e., MFT ignores the remainder of such a line.

• Five or more ‘%’ signs should not be used.

(The above description was edited from ‘mft.web’, written by D.E. Knuth.) The program accepts the following options, as well as the standard ‘-help’ and

‘-version’ (see Section 3.2 [Common options], page 7):

‘-change=chfile[.ch]’

Apply the change file chfile as with Tangle and Weave (see Chapter 8 [WEB], page 38).

‘-style=mftfile[.mft]’

Read mftfile before anything else; a MFT style file typically contains only MFT directives as described above. The default style file is named ‘plain.mft’, which defines this properly for programs using plain Metafont. The MFT files is searched along the MFTINPUTS path; see section “Supported file formats” in Kpathsea. Other examples of MFT style files are ‘cmbase.mft’, which defines formatting rules for the macros defined in ‘cm.base’, and ‘e.mft’, which was used in the production of Knuth’s Volume E, Computer Modern Typefaces. Using an appropriate MFT style file, it is also possible to configure MFT for typesetting MetaPost sources. However, MFT does not search the usual places for MetaPost input files.

If you use eight-bit characters in the input file, they are passed on verbatim to the TEX output file; it is up to you to configure TEX to print these properly.

6 MetaPost: Creating technical illustrations

MetaPost is a picture-drawing language similar to Metafont (see Chapter 5 [Metafont], page 22), but instead of outputting bitmaps in a “font”, it outputs PostScript commands. It’s primarily intended for creating technical illustrations.

MetaPost also provides for arbitrary integration of text and graphics in a natural way, using any typesetter (TEX and Troff are both supported) and a number of other subsidiary programs, described below.

6.1 mpost invocation

MetaPost (installed as mpost) reads a series of pictures specified in the MetaPost programming language, and outputs corresponding PostScript code. This section merely describes the options available in the Web2c implementation. For a complete description of the MetaPost language, see AT&T technical report CSTR-162, generally available as the file ‘texmf/doc/metapost/mpman.ps’, where texmf is the root of TEX directory structure.

See also http://cm.bell-labs.com/who/hobby/MetaPost.html.

Also, a standard MetaPost package for drawing graphs is documented in AT&T technical report CSTR-164, available as the file ‘mpgraph.ps’, generally stored alongside ‘mpman.ps’.

MetaPost processes its command line and determines its memory dump (mem) file in a way exactly analogous to Metafont and TEX (see Section 4.1 [tex invocation], page 14, and see Section 3.5.2 [Memory dumps], page 10). Synopses:

     mpost [option]... [mpname[.mp]] [mp-commands]
     mpost [option]... \first-line
     mpost [option]... &mem args

MetaPost searches the usual places for the main input file mpname (see section “Supported file formats” in Kpathsea), extending mpname with ‘.mp’ if necessary. To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

MetaPost writes its PostScript output to a series of files ‘basempname.nnn’ (or perhaps ‘basempname.ps’, very occasionally ‘basempname.tfm’), where nnn are the figure numbers specified in the input, typically to the beginfig macro, and basempname is the basename of mpname, or ‘mpout’ if no input file was specified. MetaPost uses the ‘.ps’ extension when the figure number is out of range, e.g., if you say beginfig(-1).

You can use the output files as figures in a TEX document just as with any other PostScript figures. For example, with this TEX command:

     \special{psfile="filename"}

or by using ‘epsf.tex’ (see section “EPSF macros” in Dvips).

The MetaPost construct

     btex ... tex-input ... etex

calls MakeMPX to generate a MPX file containing a MetaPost picture expression corresponding to tex-input (see Section 6.4 [makempx invocation], page 31).

The construct

     verbatimtex ... tex-input ... etex

simply passes the tex-input through to MakeMPX and thus to TEX. For example, if you are using LaTEX, your MetaPost input file must start with a verbatimtex block that gives

the necessary \documentclass (or \documentstyle) \begin{document} command. You

will also need to set the enviroment variable TEX to ‘latex’ (see Section 6.4 [makempx invocation], page 31).

tex-input need not be specifically TEX input; it could also be Troff. In that case, you will need the ‘-m pictures’ Troff macro package (unfortunately absent from many Troff implementations), or an equivalent such as the ‘-m pspic’ macros from GNU groff described in grops(1).

Other typesetters can be supported with no change to MetaPost itself; only MakeMPX needs to be updated.

Naturally, you must use fonts that are supported by the typesetter; specifically, you’ll probably want to use standard PostScript fonts with Troff. And only the TEX system understands Computer Modern or other Metafont fonts; you can also use PostScript fonts with TEX, of course.

MetaPost-generated PostScript figures which do use Computer Modern fonts for labels cannot be directly previewed or printed. Instead, you must include them in a TEX document and run the resulting DVI file through Dvips to arrange for the downloading of the required fonts (see section “Fonts in figures” in Dvips). To help with this, the MetaPost distribution provides a small TEX file ‘mproof.tex’ which is typically called as:

     tex mproof mp-output-files... ; dvips mproof -o

The resulting file ‘mproof.ps’ can then be printed or previewed.

To generate EPSF files, set the internal MetaPost variable prologues positive. To make the output files self-contained, use only standard PostScript fonts. MetaPost reads the same ‘psfonts.map’ file as Dvips, to determine PostScript fonts that need to be downloaded (see section “psfonts.map” in Dvips).

MetaPost can write output files, via the write primitive; this opens a security hole. See Section 4.1 [tex invocation], page 14.

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-kpathsea-debug=number’
‘-ini’
‘-mem=memname’
‘-progname=string’
‘-translate-file=tcxfile’

These options are common to TEX, Metafont, and MetaPost. See Section 3.2 [Common options], page 7.

‘-T’

‘-troff’ Set the prologues internal variable to 1, and use makempx -troff to generate MPX files.

‘-tex=texprogram’

When this option is given, the program texprogram is used to typeset the labels.

6.2 inimpost invocation

inimpost is the “initial” form of MetaPost, which does lengthy initializations avoided by the “virgin” (vir) form, so as to be capable of dumping ‘.mem’ files (see Section 3.5.2 [Memory dumps], page 10). For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 6.1 [mpost invocation], page 29. The only memory dump file commonly used with MetaPost is the default, ‘plain.mem’,

also known as ‘mpost.mem’ (again, see Section 3.5.2 [Memory dumps], page 10). It is created by default during installation, but you can also do so by hand if necessary (e.g., if a MetaPost update is issued):

     inimpost '\input plain dump'

(The quotes prevent interpretation of the backslashes from the shell.) Then install the

resulting ‘plain.mem’ in ‘$(memdir)’ (‘/usr/local/share/texmf/web2c’ by default), and

link ‘mpost.mem’ to it.

MetaPost also provides a mem file with all the features of plain Metafont, called ‘mfplain.mem’. You can create that in the same way; just replace ‘plain’ in the above command with ‘mfplain’. ‘mfplain.mem’ file lets you directly process Metafont source files with MetaPost, producing character proofs (one file for each character) similar to those produced with Metafont in proof mode and GFtoDVI (see Section 5.6 [gftodvi invocation], page 26).

6.3 virmpost invocation

virmpost is the “virgin” form of MetaPost, which avoids the lengthy initializations done by the “initial” (ini) form, and is thus what is generally used for production work. For a detailed comparison of virgin and initial forms, see Section 3.5.1 [Initial and virgin], page 9.

For a list of options and other information, see Section 6.1 [mpost invocation], page 29.

6.4 MakeMPX: Support MetaPost labels

In MetaPost, labels can be typeset using any document processor; the Web2c implementation supports TEX and Troff. MakeMPX translates the labels from the typesetting language back into low-level MetaPost commands in a so-called mpx file, so text can be manipulated like other graphic objects. It is invoked automatically by MetaPost. Synopsis:

     makempx [-troff] mpfile mpxfile

The input comes from mpfile (no path searching is done), and the output goes to mpxfile. However, if the file mpxfile already exists, and is newer than mpfile, then nothing is done (presumably the file is up-to-date).

Otherwise:

1. MPto is run to extract the label text from the MetaPost source file mpfile (see Section 6.7 [mpto invocation], page 34).

2. The typesetting program itself is run, either TEX or Troff (see below). If TEX, and the file named by the MPTEXPRE environment variable exists (‘mptexpre.tex’ by default), that file is prepended to the input from the MetaPost file.

3. The typesetter output (a DVI file in the case of TEX, Ditroff output for Troff) is translated back to MetaPost, by DVItoMP (see Section 6.5 [dvitomp invocation], page 33) or DMP (see Section 6.6 [dmp invocation], page 33) respectively.

If any of the above steps fail, for example if there was a typesetting mistake in the original mpfile, output may be left in files named ‘mpxerr.{log,tex,dvi}’ (TEX) or ‘mpxerr{,.t}’ (Troff), so you can diagnose the problem.

The ‘-troff’ option to MPto selects the Troff commands, rather than TEX. MetaPost supplies this automatically if the ‘-T’ or ‘-troff’ option was specified to MetaPost.

The MPX file created by MakeMPX is a sequence of MetaPost picture expressions, one for every label in the original MetaPost input file.

The names of the commands run by MakeMPX, and the directory added to the shell search PATH for the commands’ location, are overridden by environment variables. Here is a list:

MAKEMPX_BINDIR

The directory added to the PATH. Default is the ‘$(bindir)’ Make directory, which in turn is set from the configure-time ‘--bindir’, ‘--exec-prefix’ and ‘--prefix’ options; if nothing else is specified, the default is file ‘/usr/local’.

NEWER The command run to determine if mpxfile is out of date with respect to mpfile; default is ‘newer’.

MPTOTEX The command run to extract MetaPost labels in TEX format; default is ‘mpto

          -tex’.

MPTOTR Likewise, for Troff; default is ‘mpto -troff’.

DVITOMP The command run to convert TEX output back to MetaPost; default is

          ‘dvitomp’.

DMP Likewise, for Troff; default is ‘dmp’.

TEX The command run to typeset the labels in TEX; default is ‘tex’. If you use LaTEX, set this to latex, and supply an appropriate verbatimtex header in the MP source (see Section 6.1 [mpost invocation], page 29).

TROFF Likewise, for Troff; default is ‘'eqn -d\$\$ | troff -Tpost'’. You may need to replace ‘-Tpost’ by ‘-Tterm’, where term is the PostScript device name for your Troff implementation, e.g., ‘ps’ or ‘psc’; see troff(1). If you change this, you will also need to set the TRFONTS environment variable or configuration value to point to the appropriate font directory, traditionally

          ‘/usr/lib/font/devterm’.

6.5 DVItoMP: DVI to MPX conversion

DVItoMP converts DVI files into low-level MetaPost commands in a so-called MPX file. This program is generally invoked only by MakeMPX (see Section 6.4 [makempx invocation], page 31). Synopsis:

     dvitomp dvifile[.dvi] [mpxfile[.mpx]]

If mpxfile is not specified, the output goes to the basename of dvifile extended with ‘.mpx’,

e.g., ‘dvitomp /wherever/foo.dvi’ creates ‘./foo.mpx’.

The only options are ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7).

6.6 DMP: Ditroff to MPX conversion

DMP converts device-independent Troff (ditroff) output files into low-level MetaPost commands in a so-called MPX file. This program is generally invoked by MakeMPX (see Section 6.4 [makempx invocation], page 31). Synopsis:

     dmp [ditroff-file [mpxfile]]

If ditroff-file is not specified, input comes from standard input; and if mpxfile is not specified, output goes to standard output.

DMP was written to process the output of a Troff pipeline fed the output of mpto -troff (see Section 6.7 [mpto invocation], page 34). DMP understands all the ‘Dc’ graphics functions that dpost does, but it ignores ‘x X’ device control functions such as ‘x X

SetColor:...’, ‘x X BeginPath:’, and ‘x X DrawPath:...’.

The available font names are defined in the support file ‘trfonts.map’, which DMP looks for along the MPSUPPORT path.

Another support file ‘trchars.adj’, also looked for along the MPSUPPORT path, contains a character adjustment table which should reflect the shift amounts found in the standard PostScript prologue for Troff and dpost found in the TRFONTS directory. Such an adjustment table is unnecessary for some Troff implementations, in which case ‘trchars.adj’ should be replaced by an empty file—but it must still exist.

DMP was written for one particular Troff implementation, and it unfortunately has many built-in assumptions about the output and fonts file formats used by Troff, which may not be satisfied in other environments. In particular, GNU groff uses some extensions in its file formats described in groff font(5) and groff out(5) which make its output completely unusable for DMP. On the other hand, the Troff version found in Sun Solaris 2.x, and perhaps other systems derived from System V R4, works fine with the default settings.

If you run into trouble and want to adapt DMP to other systems, you might have to try the following (this is primarily for hackers):

• If DMP complains about a missing font table (e.g., ‘Cannot find TR’), your Troff may not support the device ‘post’. Check troff(1) for the devices supported by your Troff and set the TROFF environment variable appropriately (see above). Also, locate the appropriate font directory and set the TRFONTS variable as needed.

• If DMP complains about a missing font description file (e.g., ‘Font TR was not in map file’), your version of Troff may be using internal font names different from those in the distributed ‘trfonts.map’; e.g., TR and TI instead of R and I for Times-Roman and Times-Italic. In this case, you may have to adapt ‘trfonts.map’ and perhaps also ‘trchars.adj’ in the MetaPost support directory (‘texmf/metapost/support’ by default).

• If DMP still complains that it cannot parse the font description files or the Troff output (e.g., ‘TR has a bad line in its description file’, you are probably out of luck and have to hack the DMP program (in ‘web2c/mpware/dmp.c’). Such problems may be caused by subtle differences in the file formats, such as use of tabs vs. spaces as field separators or decimal vs. octal vs. hex format for font metric data. A reasonably good description of the expected Troff file formats can be found in AT&T technical report CSTR-54 (Troff User’s Manual, Revised 1992). Documentation on the subtle differences in other Troff implementation is harder to find except for GNU groff, where it’s all documented in the above-mentioned groff font(5) and groff out(5). Any contributions to improve the portability of DMP or to make it work with GNU groff are welcome, of course.

(Some of the above description was edited from the ‘dmp.c’ source file, written by John Hobby.)

The only options are ‘--help’ and ‘--version’ (see Section 3.2 [Common options], page 7).

6.7 MPto: Extract labels from MetaPost input

MPto extracts the labels from a MetaPost input file; this is the contents of any btex...etex and verbatimtex...etex sections. This program is generally invoked by MakeMPX (see Section 6.4 [makempx invocation], page 31). Synopsis:

 mpto [option]... mpfile

The input comes from mpfile; no path searching is done. The output goes to standard output. Leading and trailing spaces and tabs are removed, and various predefined typesetter commands are included at the beginning of and end of the file and of each section.

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-troff’ Surround the MetaPost sections with Troff commands.

‘-tex’ Surround the MetaPost sections with TEX commands. This is the default.

6.8 Newer: Compare file modification times

Newer compares file modification times. Synopsis:

     newer src dependent

Newer exits successfully if the file src exists and is older as dependent, i.e., the modification time (mtime) of src is greater than that of dependent. See section “Attribute Meanings” in GNU C Library.

Although this could be written as a Perl script (see section “File Operations” in Perl) or using the ‘--full-time’ option supported by ls (see section “ls invocation” in GNU file utilities), it seems undesirable to depend on such independent, and sadly non-universal, programs.

This is used by MakeMPX (see Section 6.4 [makempx invocation], page 31).

7 BibTEX: Bibliographies

BibTEX automates much of the job of typesetting bibliographies, and makes bibliography entries reusable in many different contexts.

7.1 BibTEX invocation

BibTEX creates a printable bibliography (‘.bbl’) file from references in a ‘.aux’ file, generally written by TEX or LaTEX. The ‘.bbl’ file is then incorporated on a subsequent run. The basic bibliographic information comes from ‘.bib’ files, and a BibTEX style (‘.bst’) file controls the precise contents of the ‘.bbl’ file. Synopsis:

     bibtex [option]... auxfile[.aux]

The output goes to the basename of auxfile extended with ‘.bbl’; for example, ‘bibtex /wherever/foo.aux’ creates ‘./foo.bbl’. BibTEX also writes a log file to the basename of auxfile extended with ‘.blg’.

The names of the ‘.bib’ and ‘.bst’ files are specified in the ‘.aux’ file as well, via the ‘\bibliography’ and ‘\bibliographystyle’ (La)TEX macros. BibTEX searches for ‘.bib’ files using the BIBINPUTS and TEXBIB paths, and for ‘.bst’ files using BSTINPUTS (see section “Supported file formats” in Kpathsea). It does no path searching for ‘.aux’ files.

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-terse’ Suppress the program banner and progress reports normally output.

‘-min-crossrefs=n’

If at least n (2 by default) bibliography entries refer to another entry e via their crossref field, include e in the .bbl file, even if it was not explicitly referenced in the .aux file. For example, e might be a conference proceedings as a whole, with the cross-referencing entries being individual articles published in the proceedings. In some circumstances, you may want to avoid these automatic inclusions altogether; to do this, make n a sufficiently large number.

See also:

‘btxdoc.tex’

Basic LaTEXable documentation for general BibTEX users.

‘btxhak.tex’

LaTEXable documentation for style designers.

‘btxdoc.bib’

BibTEX database file for the two above documents.

‘xampl.bib’

Example database file with all the standard entry types.

‘ftp://ftp.math.utah.edu/pub/tex/bib/’

A very large ‘.bib’ and ‘.bst’ collection, including references for all the standard TEX books and a complete bibliography for TUGboat.

7.2 Basic BibTEX style files

Here are descriptions of the four standard and four semi-standard basic BibTEX styles. ‘CTAN:/biblio/bibtex’ contains these and many more (for CTAN info, see section “unixtex.ftp” in Kpathsea).

plain Sorts entries alphabetically, with numeric labels. Generally formatted according to van Leunen’s A Handbook for Scholars. The other style files listed here are based on plain.

abbrv First names, month names, and journal names are abbreviated.

acm Names are printed in small caps.

alpha Alphanumeric labels, e.g., ‘Knu66’.

apalike No labels at all; instead, the year appears in parentheses after the author. Should be used in conjunction with ‘apalike.tex’ (plain TEX) or ‘apalike.sty’ (LaTEX), which also changes the citations in the text to be

          ‘(author, year)’.

ieeetr Numeric labels, entries in citation order, ieee abbreviations, article titles in quotes.

siam Numeric labels, alphabetic order, Math. Reviews abbreviations, names in small caps.

unsrt Lists entries in citation order, i.e., unsorted.

btxbst.doc

The template file and documentation for the standard styles.

8 WEB: Literate programming

WEB languages allow you to write a single source file that can produce both a compilable program and a well-formatted document describing the program in as much detail as you wish to prepare. Writing in this kind of dual-purpose language is called literate programming. (The Usenet newsgroup ‘comp.programming.literate’ and the mailing list litprog@shsu.edu are devoted to this subject; they are gatewayed to each other.)

WEB-like languages have been implemented with many pairs of base languages: Cweb provides C and Troff (see Appendix B [References], page 55); CWEB provides C and TEX (‘CTAN:/web/c_cpp/cweb’); Spiderweb provides C, C++, Awk, Ada, many others, and TEX (‘CTAN:/web/spiderweb’); and, of course, the original WEB provides Pascal and TEX, the implementation languages for the original TEX, Metafont, MetaPost, and related programs to come from the TEX project at Stanford.

The original WEB language is documented in the file ‘webman.tex’, which is included in the ftp://ftp.tug.org/tex/lib.tar.gz archive (and available in many other places, of course).

8.1 Tangle: Translate WEB to Pascal

Tangle creates a compilable Pascal program from a WEB source file (see Chapter 8 [WEB], page 38). Synopsis:

     tangle [option]... webfile[.web] [changefile[.ch]]

The Pascal output is written to the basename of webfile extended with ‘.p’; for example, ‘tangle /wherever/foo.web’ creates ‘./foo.p’. Tangle applies changefile to webfile before writing the output; by default, there is no change file.

If the program makes use of the WEB string facility, Tangle writes the string pool to the basename of webfile extended with ‘.pool’.

The Pascal output is packed into lines of 72 characters or less, with the only concession to readability being the termination of lines at semicolons when this can be done conveniently.

The program accepts the following options, as well as the standard ‘--help’ and ‘--version’ (see Section 3.2 [Common options], page 7):

‘-length=number’

The number of characters that are considered significant in an identifier. Whether underline characters are counted depends on the ‘-underline’ option. The default value is 32, the original tangle used 7, but this proved too restrictive for use by Web2c.

‘-lowercase’
‘-mixedcase’
‘-uppercase’

These options specify the case of identifiers in the output of tangle. If ‘-uppercase’ (‘-lowercase’) is specified, tangle will convert all identfiers to uppercase (lowercase). The default is ‘-mixedcase’, which specifies that the case will not be changed.

‘-underline’

When this option is given, tangle does not strip underline characters from identifiers.

‘-loose’

‘-strict’ These options specify how strict tangle must be when checking identifiers for equality. The default is ‘-loose’, which means that tangle will follow the rules set by the case-smashing and underline options above. If ‘-strict’ is set, then identifiers will always be stripped of underlines and converted to uppercase before checking whether they collide.

8.2 Weave: Translate WEB to TEX

Weave creates a TEX document from a WEB source file (see Chapter 8 [WEB], page 38), assuming various macros defined in ‘webmac.tex’. It takes care of typographic details such as page layout, indentation, and italicizing identifiers. It also automatically gathers and outputs extensive cross-reference information. Synopsis:

     weave [option]... webfile[.web] [changefile[.ch]]

The output is to the basename of webfile extended with ‘.tex’; for example, ‘weave /wherever/foo.web’ creates ‘./foo.tex’. Weave applies changefile to webfile before writing the output; by default, there is no change file.

The program accepts the following option, as well as the standard ‘-verbose’, ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-x’ Omit the cross-reference information: the index, the list of WEB module names, and the table of contents (an empty ‘CONTENTS.tex’ file will still be written when the Weave output file is processed by TEX using the default ‘webmac.tex’, though).

Conventionally, WEB programmers should define the TEX \title macro at the beginning of the source file. Also, to get output of only changed modules, one can say \let\maybe=\iffalse (usually as the first change in the change file).

8.3 Pooltype: Display WEB pool files

Pooltype shows the so-called string number of each string in a WEB pool file (see Chapter 8 [WEB], page 38), as output by Tangle (see Section 8.1 [tangle invocation], page 38), including the first 256 strings corresponding to the possible input characters. Pooltype primarily serves as an example of WEB conventions to implementors of the TEX system. Synopsis:

     pooltype [option]... poolfile[.pool]

No path searching is done for poolfile. Output is to standard output.

The only options are ‘--help’ and ‘--version’ (see Section 3.2 [Common options], page 7).

As an example of the output, here is the (edited) output for ‘tex.pool’:

0: "^^@"
1: "^^A"
...
255: "^^ff"
256: "pool size"
...
1314: "Using character substitution: "
(23617 characters in all.)

In Metafont and MetaPost, the first 256 characters are actually represented as single bytes (i.e., themselves), not in the ‘^^’ notation. Consider Pooltype as showing the results after conversion for output.

9 DVI utilities

TEX outputs a file in DVI (DeVice Independent) format as a compact representation of the original document. DVI files can be translated to meet the requirements of a real physical device, such as PostScript printers (see section “Introduction” in Dvips), PCL printers (see dvilj(1)), and X displays (see xdvi(1)). In fact, DVI translators are available for virtually all common devices: see ‘CTAN:/dviware’ (for CTAN info, see section “unixtex.ftp” in Kpathsea).

For the precise definition of the DVI file format, see (for example) the source file

‘web2c/dvitype.web’.

The DVI-processing programs in the Web2c distribution are not device drivers; they perform generic utility functions.

9.1 DVIcopy: Canonicalize virtual font references

DVIcopy reads a DVI file, expands any references to virtual fonts (see section “Virtual fonts” in Dvips) to base fonts, and writes the resulting DVI file. Thus you can use virtual fonts even if your DVI processor does not support them, by passing the documents through DVIcopy first. Synopsis:

     dvicopy [option]... [indvi[.dvi] [outdvi[.dvi]]]

DVIcopy reads standard input if indvi is not specified, and writes standard output if outdvi is not specified.

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-magnification=integer’

Override existing magnification in indvi with integer; 1000 specifies no magnification. This is equivalent to setting TEX’s \mag parameter.

‘-max-pages=n’

Process n pages; default is one million.

‘-page-start=page-spec’

Start at the first page matching page-spec, which is one or more (signed) integers separated by periods, corresponding to TEX’s \count0...9 parameters at \shipout time; ‘*’ matches anything. Examples: ‘3’, ‘1.*.-4’.

9.2 DVItype: Plain text transliteration of DVI files

DVItype translates a DeVice Independent (DVI) file (as output by TEX, for example) to a plain text file that humans can read. It also serves as a DVI-validating program, i.e., if DVItype can read a file, it’s correct. Synopsis:

     dvitype [option]... dvifile[.dvi]

DVItype does not read any bitmap files, but it does read TFM files for fonts referenced in dvifile. The usual places are searched (see section “Supported file formats” in Kpathsea).

To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

Output goes to standard output.

The program accepts the following options, as well as the standard ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-dpi=real’

Do pixel movement calculations at real pixels per inch; default 300.0.

‘-magnification=integer’

Override existing magnification in indvi with integer; 1000 specifies no magnification. This is equivalent to setting TEX’s \mag parameter.

‘-max-pages=n’

Process n pages; default is one million.

‘-output-level=n’

Verbosity level of output, from 0 to 4 (default 4): • 0: Global document information only.

• 1: Most DVI commands included, and typeset characters summarized.

• 2: Character and movement commands explicitly included.

• 3: DVI stack and current position calculations included.

• 4: Same information as level 3, but DVItype does random positioning in the file, reading the DVI postamble first.

‘-page-start=page-spec’

Start at the first page matching page-spec, which is one or more (signed) integers separated by periods, corresponding to TEX’s \count0...9 parameters at \shipout time; ‘*’ matches anything. Examples: ‘1’, ‘5.*.-9’.

‘-show-opcodes’

Show numeric opcode values (in decimal) for DVI commands, in braces after the command name. This can help in debugging DVI utilities. We use decimal because in the DVI format documentation (in ‘dvitype.web’, among others) the opcodes are shown in decimal.

9.2.1 DVItype output example

As an example of the output from DVItype (see section above), here is its (abridged) translation of the ‘story.dvi’ resulting from running the example in The TEXbook, with

‘-output-level=4’ and ‘-show-opcodes’ on.
     ...
     Options selected:
       Starting page = *
       Maximum number of pages = 1000000
       Output level = 4 (the works)
       Resolution = 300.00000000 pixels per inch
     numerator/denominator=25400000/473628672

magnification=1000;       0.00006334 pixels per DVI unit
' TeX output 1992.05.17:0844'
Postamble starts at byte 564.
maxv=43725786, maxh=30785863, maxstackdepth=3, totalpages=1
Font 33: cmsl10---loaded at size 655360 DVI units
Font 23: cmbx10---loaded at size 655360 DVI units
Font 0: cmr10---loaded at size 655360 DVI units
42: beginning of page 1
87: push {141}
level 0:(h=0,v=0,w=0,x=0,y=0,z=0,hh=0,vv=0)
88: down3 -917504 {159} v:=0-917504=-917504, vv:=-58
92: pop {142}
...
104: putrule {137} height 26214, width 30785863 (2x1950 pixels)
113: down3 5185936 {159} v:=655360+5185936=5841296, vv:=370
117: push {141}
level 1:(h=0,v=5841296,w=0,x=0,y=0,z=0,hh=0,vv=370)
118: right4 12265425 {146} h:=0+12265425=12265425, hh:=777
[ ]
123: fntdef1 23 {243}: cmbx10
145: fntnum23 {194} current font is cmbx10
146: setchar65 h:=12265425+569796=12835221, hh:=813
147: w3 251220 {150} h:=12835221+251220=13086441, hh:=829
151: setchar83 h:=13086441+418700=13505141, hh:=856
...
164: setchar82 h:=17448202+565245=18013447, hh:=1142
165: x0 -62805 {152} h:=18013447-62805=17950642, hh:=1138
166: setchar89 h:=17950642+569796=18520438, hh:=1174
[A SHORT STORY]
167: pop {142}
level 1:(h=0,v=5841296,w=0,x=0,y=0,z=0,hh=0,vv=370)
...
550: pop {142}
level 0:(h=0,v=42152922,w=0,x=0,y=0,z=0,hh=0,vv=2670)
551: down3 1572864 {159} v:=42152922+1572864=43725786, vv:=2770
555: push {141}
level 0:(h=0,v=43725786,w=0,x=0,y=0,z=0,hh=0,vv=2770)
556: right4 15229091 {146} h:=0+15229091=15229091, hh:=965
561: setchar49 h:=15229091+327681=15556772, hh:=986
[ 1]
562: pop {142}
level 0:(h=0,v=43725786,w=0,x=0,y=0,z=0,hh=0,vv=2770)
563: eop {140}

Explanation:

• The DVItype options are recorded at the beginning, followed by global information

about the document, including fonts used.

• Each DVI command is preceded by its byte position in the file (‘42:’, ‘87:’, …), and (because of the ‘-show-opcodes’) followed by its decimal opcode value in braces (‘{141}’, ‘{142}’, …).

• The ‘level’ lines record information about the DVI stack; ‘h’ and ‘v’ define the current position in DVI units, while ‘hh’ and ‘vv’ are the same in pixels.

• Text sequences are summarized in brackets, as in ‘[A SHORT STORY]’ and the ‘[ 1]’.

10 Font utilities

The Web2c programs described here convert between various TEX-related font formats; the first section below briefly describes the formats. GFtoPK is the only one that is routinely used, as Metafont outputs GF format, but it’s most efficient for device drivers to use PK.

The precise definitions of the PK, GF, TFM, PL, VF, and VPL formats mentioned below are in the source files that read them; ‘pktype.web’, ‘gftype.web’, ‘tftopl.web’, etc.

10.1 Font file formats

(For another perspective on this, see section “Font concepts” in Dvips).

Font files come in several varieties, with suffixes like:

     .tfm  .*pk  .*gf  .*pxl (obsolete) .pl  .mf  .vf  .vpl

Each represents a file format.

A TFM (TEX font metric) file is a compact binary file that contains information about each character in a font, about combinations of characters within that font, and about the font as a whole. The font metric information contained in TFM files is device-independent units is used by TEX to do typesetting. Unlike the bitmap (raster) fonts described below, TFM font files contain no information about the shapes of characters. They describe rectangular areas and combinations thereof, but not what will eventually be printed in those areas.

Since TEX does scaling calculations, one TFM file serves for all magnifications of a given typeface. On the other hand, the best printed results are obtained when magnified (or reduced fonts) are not produced geometrically (as done by PostScript, for example) but rather optically, with each size a separate design (as done with Computer Modern and the EC fonts, for example); then a separate TFM file is needed for each size.

At any rate, TEX produces a DVI (DeVice Independent) file from your source document. In order to print DVI files on real devices, you need font files defining digitized character shapes and other data. Then previewers and printer-driver programs can translate your DVI files into something usable by your monitor or printer. Bitmap fonts come with suffixes such as ‘.600pk’ or ‘.600gf’ or ‘.3000pxl’, where the ‘600’ is the horizontal dots-per-inch resolution at which the font was produced, and the ‘pk’ or ‘gf’ or ‘pxl’ indicates the font format. Outline fonts in PostScript Type 1 format have suffixes such as ‘.pfa’ or ‘.pfb’.

Fonts in pk (packed) format are in the tightly packed raster format that is pretty much the standard today. They take up less space than fonts in the gf (generic font) format that Metafont generates, and far less space than fonts in pxl format. Fonts in pxl format take up gross amounts of disk space and permit only 128 characters. They are obsolete.

Font files with the ‘.pl’ (property list) suffix are the plain text (human-readable) analog of the binary ‘.tfm’ files. The TFtoPL and PLtoTF programs convert between the two formats (see Section 10.6 [tftopl invocation], page 50 and Section 10.7 [pltotf invocation], page 52).

Font files with the ‘.mf’ suffix are in Metafont source format. These are the files used by Metafont to generate rastered fonts for specific typefaces at specific magnifications for the specific resolution and type of mapping used by your device.

The suffix ‘.vf’ identifies “virtual font” files, for which ‘.vpl’ is the human-readable analog. See See Section 10.8 [vftovp invocation], page 52, and Section 10.9 [vptovf invocation], page 53. For further discussion of virtual fonts, see ‘CTAN:/doc/virtual-fonts.knuth’, ‘CTAN:/help/virtualfonts.txt’, and section “Virtual fonts” in Dvips.

(This section is based on documentation in the original Unix TEX distribution by Pierre MacKay and Elizabeth Tachikawa.)

10.2 GFtoPK: Generic to packed font conversion

GFtoPK converts a generic font (GF) file output by, for example, Metafont (see Section 5.1 [mf invocation], page 22) to a packed font (PK) file. PK files are considerably smaller than the corresponding gf files, so they are generally the bitmap font format of choice. Some DVI-processing programs, notably Dvips, only support PK files and not GF files. Synopsis:

     gftopk [option]... gfname.dpi[gf] [pkfile]

The font gfname is searched for in the usual places (see section “Glyph lookup” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

The suffix ‘gf’ is supplied if not already present. This suffix is not an extension; no ‘.’ precedes it: for instance, ‘cmr10.600gf’.

If pkfile is not specified, the output is written to the basename of ‘gfname.dpipk’, e.g.,

‘gftopk /wherever/cmr10.600gf’ creates ‘./cmr10.600pk’.

The only options are ‘--verbose’, ‘--help’, and ‘--version’ (see Section 3.2 [Common options], page 7).

10.3 PKtoGF: Packed to generic font conversion

PKtoGF converts a packed font (PK) file to a generic font (GF) file. Since PK format is much more compact than GF format, the most likely reason to do this is to run GFtype (see Section 10.5 [gftype invocation], page 48) on the result, so you can see the bitmap images. Also, a few old utility programs do not support PK format. Synopsis:

     pktogf [option]... pkname.dpi[pk] [gffile]

The font pkname is searched for in the usual places (see section “Glyph lookup” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

The suffix ‘pk’ is supplied if not already present. This suffix is not an extension; no ‘.’ precedes it: for instance, ‘cmr10.600pk’.

If gffile is not specified, the output is written to the basename of ‘pkname.dpigf’, e.g.,

‘pktogf /wherever/cmr10.600pk’ creates ‘./cmr10.600gf’.

The only options are ‘--verbose’, ‘--help’, and ‘--version’ (see Section 3.2 [Common options], page 7).

10.4 PKtype: Plain text transliteration of packed fonts

PKtype translates a packed font (PK) bitmap file (as output by GFtoPK, for example) to a plain text file that humans can read. It also serves as a PK-validating program, i.e., if PKtype can read a file, it’s correct. Synopsis:

     pktype pkname.dpi[pk]

The font pkname is searched for in the usual places (see section “Glyph lookup” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

The suffix ‘pk’ is supplied if not already present. This suffix is not an extension; no ‘.’ precedes it: for instance, ‘cmr10.600pk’.

The translation is written to standard output.

The only options are ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7). As an example of the output, here is the (abridged) translation of the letter ‘K’ in

‘cmr10’, as rendered at 600 dpi with the mode ‘ljfour’ from modes.mf (available from

‘ftp://ftp.tug.org/tex/modes.mf’).
      955:  Flag byte = 184  Character = 75  Packet length = 174
        Dynamic packing variable = 11
        TFM width = 815562  dx = 4259840
        Height = 57  Width = 57  X-offset = -3  Y-offset = 56
        [2]23(16)17(8)9(25)11(13)7(27)7(16)7(28)4(18)7(28)2(20)7(27)...
        ...
        (14)9(24)12(5)[2]23(13)21

Explanation:

‘955’ The byte position in the file where this character starts.

‘Flag byte’
‘Dynamic packing variable’

Related to the packing for this character; see the source code.

‘Character’

The character code, in decimal.

‘Packet length’

The total length of this character definition, in bytes.

‘TFM width’

The device-independent (TFM) width of this character. It is 2^24 times the ratio of the true width to the font’s design size.

‘dx’ The device-dependent width, in scaled pixels, i.e., units of horizontal pixels times 2^16.

‘Height’

‘Width’ The bitmap height and width, in pixels.

‘X-offset’
‘Y-offset’

Horizontal and vertical offset from the upper left pixel to the reference (origin) pixel for this character, in pixels (right and down are positive). The reference pixel is the pixel that occupies the unit square in Metafont; the Metafont reference point is the lower left hand corner of this pixel. Put another way, the x-offset is the negative of the left side bearing; the right side bearing is the horizontal escapement minus the bitmap width plus the x-offset.

‘[2]23(16)...’

Finally, run lengths of black pixels alternate with parenthesized run lengths of white pixels, and brackets indicate a repeated row.

10.5 GFtype: Plain text transliteration of generic fonts

GFtype translates a generic font (GF) bitmap file (as output by Metafont, for example) to a plain text file that humans can read. It also serves as a GF-validating program, i.e., if GFtype can read a file, it’s correct. Synopsis:

gftype [option]... gfname.dpi[gf]

The font gfname is searched for in the usual places (see section “Glyph lookup” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

The suffix ‘gf’ is supplied if not already present. This suffix is not an extension; no ‘.’ precedes it: for instance, ‘cmr10.600gf’.

The translation is written to standard output. The program accepts the following options, as well as the standard ‘-help’ and

‘-version’ (see Section 3.2 [Common options], page 7):

‘-images’ Show the characters’ bitmaps using asterisks and spaces.

‘-mnemonics’

Translate all commands in the GF file.

As an example of the output, here is the (abrdiged) translation of the letter ‘K’ in ‘cmr10’, as rendered at 600 dpi with the mode ‘ljfour’ from ‘modes.mf’ (available from

ftp://ftp.tug.org/tex/modes.mf), with both ‘-mnemonics’ and ‘-images’ enabled.

GFtype outputs the information about a character in two places: a main definition and a one-line summary at the end. We show both. Here is the main definition:

2033: beginning of char 75: 3<=m<=60 0<=n<=56
(initially n=56) paint (0)24(12)20
2043: newrow 0 (n=55) paint 24(12)20
2047: newrow 0 (n=54) paint 24(12)20
2051: newrow 0 (n=53) paint 24(12)20
2055: newrow 7 (n=52) paint 10(21)13
2059: newrow 8 (n=51) paint 8(23)9
...
2249: newrow 8 (n=5) paint 8(23)11
2253: newrow 7 (n=4) paint 10(22)12
2257: newrow 0 (n=3) paint 24(11)22
2261: newrow 0 (n=2) paint 24(11)22
2265: newrow 0 (n=1) paint 24(11)22

2269: newrow 0 (n=0) paint 24(11)22
2273: eoc
.<--This pixel's lower left corner is at (3,57) in METAFONT coordinates
************************            ********************
************************            ********************
************************            ********************
************************            ********************
       **********                     *************
        ********                       *********
...
        ********                       ***********
       **********                      ************
************************           **********************
************************           **********************
************************           **********************
************************           **********************
.<--This pixel's upper left corner is at (3,0) in METAFONT coordinates

Explanation:

‘2033’
‘2043’

‘...’ The byte position in the file where each GF command starts.

‘beginning of char 75’

The character code, in decimal.

‘3<=m<=60 0<=n<=56’

The character’s bitmap lies between 3 and 60 (inclusive) horizontally, and between 0 and 56 (inclusive) vertically. (m is a column position and n is a row position.) Thus, 3 is the left side bearing. The right side bearing is the horizontal escapement (given below) minus the maximum m.

‘(initially n=56) paint (0)24(12)20’

The first row of pixels: 0 white pixels, 24 black pixels, 12 white pixels, etc.

‘newrow 0 (n=55) paint 24(12)20’

The second row of pixels, with zero leading white pixels on the row.

‘eoc’ The end of the main character definition.

Here is the GF postamble information that GFtype outputs at the end:

Character 75: dx 4259840 (65), width 815562 (64.57289), loc 2033

Explanation:

‘dx’ The device-dependent width, in scaled pixels, i.e., units of horizontal pixels times 2^16. The ‘(65)’ is simply the same number rounded. If the vertical escapement is nonzero, it would appear here as a ‘dy’ value.

‘width’ The device-independent (TFM) width of this character. It is 2^24 times the ratio of the true width to the font’s design size. The ‘64.57289’ is the same number converted to pixels.

‘loc’ The byte position in the file where this character starts.

10.6 TFtoPL: TEX font metric to property list conversion

TFtoPL translates a TEX font metric (TFM, see section “Metric files” in Dvips) file (as output by Metafont, for example) to property list format (a list of parenthesized items describing the font) that humans can edit or read. This program is mostly used by people debugging TEX implementations, writing font utilities, etc. Synopsis:

     tftopl [option]... tfmname[.tfm] [plfile[.pl]]

The font tfmname (extended with ‘.tfm’ if necessary) is searched for in the usual places (see section “Supported file formats” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program.

If plfile (which is extended with ‘.pl’ if necessary) is not specified, the property list file is written to standard output. The property list file can be converted back to TFM format by the companion program TFtoPL (see the next section).

The program accepts the following option, as well as the standard ‘-verbose’, ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-charcode-format=type’

Output character codes in the PL file according to type: either ‘octal’ or ‘ascii’. Default is ‘ascii’ for letters and digits, octal for all other characters. Exception: if the font’s coding scheme starts with ‘TeX math sy’ or ‘TeX math ex’, all character codes are output in octal. In ‘ascii’ format, character codes that correspond to graphic characters, except for left and right parentheses, are output as a ‘C’ followed by the single character: ‘C K’, for example. In octal format, character codes are output as the letter ‘O’ followed by octal digits, as in ‘O 113’ for ‘K’. ‘octal’ format is useful for symbol and other non-alphabetic fonts, where using ASCII characters for the character codes is merely confusing.

As an example of the output, here is the (abridged) property list translation of

‘cmr10.tfm’:
     (FAMILY CMR)
     (FACE O 352)
     (CODINGSCHEME TEX TEXT)
     (DESIGNSIZE R 10.0)
     (COMMENT DESIGNSIZE IS IN POINTS)
     (COMMENT OTHER SIZES ARE MULTIPLES OF DESIGNSIZE)
     (CHECKSUM O 11374260171)
     (FONTDIMEN
        (SLANT R 0.0)
        (SPACE R 0.333334)
        (STRETCH R 0.166667)
        (SHRINK R 0.111112)
        (XHEIGHT R 0.430555)
        (QUAD R 1.000003)
        (EXTRASPACE R 0.111112)
        )

(LIGTABLE
...
(LABEL C f)
(LIG C i O 14)
(LIG C f O 13)
(LIG C l O 15)
(KRN O 47 R 0.077779)
(KRN O 77 R 0.077779)
(KRN O 41 R 0.077779)
(KRN O 51 R 0.077779)
(KRN O 135 R 0.077779)
(STOP)
...
)
...
(CHARACTER C f
(CHARWD R 0.305557)
(CHARHT R 0.694445)
(CHARIC R 0.077779)
(COMMENT
   (LIG C i O 14)
   (LIG C f O 13)
   (LIG C l O 15)
   (KRN O 47 R 0.077779)
   (KRN O 77 R 0.077779)
   ...
   )
)
...

As you can see, the general format is a list of parenthesized properties, nested where necessary.

• The first few items (FAMILY, FACE, and so on) are the so-called headerbyte information from Metafont, giving general information about the font.

• The FONTDIMEN property defines the TEX \fontdimen values.

• The LIGTABLE property defines the ligature and kerning table. LIG properties define ligatures: in the example above, an ‘f’ (in the ‘LABEL’) followed by an ‘i’ is a ligature, i.e., a typesetting program like TEX replaces those two consecutive characters by the character at position octal ’014 in the current font—presumably the ‘fi’ ligature. KRN properties define kerns: if an ‘f’ is followed by character octal ’047 (an apostrophe), TEX inserts a small amount of space between them: 0.077779 times the design size the font was loaded at (about three-quarters of a printer’s point by default in this case, or .001 inches).

• The CHARACTER property defines the dimensions of a character: its width, height, depth, and italic correction, also in design-size units, as explained in the previous item. For our example ‘f’, the depth is zero, so that property is omitted. TFtoPL also inserts any kerns and ligatures for this character as a comment.

10.7 PLtoTF: Property list to TEX font metric conversion

PLtoTF translates a property list file (as output by TFtoPL, for example) to TEX font metric (TFM, see section “Metric files” in Dvips) format. It’s much easier for both programs and humans to create the (plain text) property list files and let PLtoTF take care of creating the binary TFM equivalent than to output TFM files directly. Synopsis:

     pltotf [option]... plfile[.pl] [tfmfile[.tfm]]

If tfmfile (extended with ‘.tfm’ if necessary) is not specified, the TFM file is written to

the basename of ‘plfile.tfm’, e.g., ‘pltotf /wherever/cmr10.pl’ creates ‘./cmr10.tfm’.

(Since TFM files are binary, writing to standard output by default is undesirable.)

The only options are ‘-verbose’, ‘-help’, and ‘-version’ (see Section 3.2 [Common options], page 7).

For an example of property list format, see the previous section.

10.8 VFtoVP: Virtual font to virtual property lists

VFtoVP translates a virtual font metric (VF, see section “Virtual fonts” in Dvips) file and its accompanying TEX font metric (TFM, see section “Metric files” in Dvips) file (as output by VPtoVF, for example) to virtual property list format (a list of parenthesized items describing the virtual font) that humans can edit or read. This program is mostly used by people debugging virtual font utilities. Synopsis:

     vftovp [option]... vfname[.vf] [tfmname[.tfm] [vplfile[.vpl]]]

The fonts vfname and tfmname (extended with ‘.vf’ and ‘.tfm’ if necessary) are searched for in the usual places (see section “Supported file formats” in Kpathsea). To see all the relevant paths, set the environment variable KPATHSEA_DEBUG to ‘-1’ before running the program. If tfmname is not specified, vfname (without a trailing ‘.vf’) is used.

If vplfile (extended with ‘.vpl’ if necessary) is not specified, the property list file is written to standard output. The property list file can be converted back to VF and TFM format by the companion program VFtoVP (see the next section).

The program accepts the following option, as well as the standard ‘-verbose’, ‘-help’ and ‘-version’ (see Section 3.2 [Common options], page 7):

‘-charcode-format=type’

Output character codes in the PL file according to type: either ‘octal’ or ‘ascii’. Default is ‘ascii’ for letters and digits, octal for all other characters. Exception: if the font’s coding scheme starts with ‘TeX math sy’ or ‘TeX math ex’, all character codes are output in octal. In ‘ascii’ format, character codes that correspond to graphic characters, except for left and right parentheses, are output as a ‘C’ followed by the single character: ‘C K’, for example. In octal format, character codes are output as the letter ‘O’ followed by octal digits, as in ‘O 113’ for ‘K’. ‘octal’ format is useful for symbol and other non-alphabetic fonts, where using ASCII characters for the character codes is merely confusing.

10.9 VPtoVF: Virtual property lists to virtual font

VPtoVF translates a virtual property list file (as output by VFtoVP, for example) to virtual font (VF, see section “Virtual fonts” in Dvips) and TEX font metric (TFM, see section “Metric files” in Dvips) files. It’s much easier for both programs and humans to create the (plain text) property list files and let VPtoVF take care of creating the binary VF and TFM equivalents than to output them directly. Synopsis:

     vptovf [option]... vplfile[.vpl] [vffile[.vf] [tfmfile[.tfm]]]

If vffile (extended with ‘.vf’ if necessary) is not specified, the VF file is written to the basename of ‘vplfile.vf’; similarly for tfmfile. For example, ‘vptovf

/wherever/ptmr.vpl’ creates ‘./ptmr.vf’ and ‘./ptmr.tfm’.

The only options are ‘-verbose’, ‘-help’, and ‘-version’ (see Section 3.2 [Common options], page 7).

10.10 Font utilities available elsewhere

The Web2c complement of font utilities merely implements a few basic conversions. Many other more sophisticated font utilities exist; most are in ‘CTAN:/fonts/utilities’ (for CTAN info, see section “unixtex.ftp” in Kpathsea). Here are some of the most commonly-requested items:

• AFM (Adobe font metric) to TFM conversion: see section “Invoking afm2tfm” in

    Dvips, and ‘CTAN:/fonts/utilities/afmtopl’.

• BDF (the X bitmap format) conversion: ftp://ftp.tug.org/tex/bdf.tar.gz.

• Editing of bitmap fonts: Xbfe from the GNU font utilities mentioned below; the

    X BDF-editing programs available from ftp://ftp.x.org/R5contrib/xfed.tar.Z
    and ftp://ftp.x.org/R5contrib/xfedor.tar.Z; and finally, if your fonts have only

128 characters, you can use the old gftopxl, pxtoch, and chtopx programs from

    ftp://ftp.tug.org/tex/web.

• PK bitmaps from PostScript fonts: gsftopk from the ‘xdvik’ distribution

    or from ‘CTAN:/fonts/utilities/gsftopk’; alternatively, ps2pk, from
    ‘CTAN:/fonts/utilities/ps2pk’.

• PostScript Type 1 font format conversion (i.e., between PFA and PFB formats):

    ftp://ftp.tug.org/tex/t1utils.tar.gz.

• Scanned image conversion: the (aging) GNU font utilities convert type specimen im-

    ages to Metafont, PostScript, etc.: ftp://prep.ai.mit.edu/pub/gnu/fontutils-
    0.6.tar.gz.
  • Virtual font creation: ‘CTAN:/fonts/utilities/fontinst’.

Appendix A Legalisms

In general, each file has its own copyright notice stating the copying permissions for that file. Following is a summary.

The Web2c system itself and most of the original WEB source files are public domain. ‘tex.web’, the MLTEX code, ‘mf.web’, and ‘bibtex.web’, are copyrighted by their au-

thors. They may be copied verbatim, but may be modified only through a ‘.ch’ file.

MetaPost-related files, including ‘mp.web’ itself, are copyrighted under X-like terms; the precise notice is included below.

Finally, almost all of the Kpathsea library is covered by the GNU Library General Public License, but part of one file is covered by the regular GNU General Public License (see section “Introduction” in Kpathsea). Therefore, the binaries resulting from a standard Web2c compilation are also covered by the GPL; so if you (re)distribute the binaries, you must also (offer to) distribute the complete source that went into those binaries. See the files ‘COPYING’ and ‘COPYING.LIB’ for complete details on the GPL and LGPL.

The following notice must be included by the terms of the MetaPost copyright. Permission to use, copy, modify, and distribute this software and its documentation for any purpose and without fee is hereby granted, provided that the above copyright notice appear in all copies and that both that the copyright notice and this permission notice and warranty disclaimer appear in supporting documentation, and that the names of AT&T Bell Laboratories or any of its entities not be used in advertising or publicity pertaining to distribution of the software without specific, written prior permission. AT&T disclaims all warranties with regard to this software, including all implied warranties of merchantability and fitness. In no event shall AT&T be liable for any special, indirect or consequential damages or any damages whatsoever resulting from loss of use, data or profits, whether in an action of contract, negligence or other tortious action, arising out of or in connection with the use or performance of this software.

Appendix B References

1. Kpathsea: See section “Top” in Kpathsea.

2. Dvips and Afm2tfm: See section “Top” in Dvips.

3. For a bibliography of formal articles and technical reports on the TEX project, see the books TEX: The Program or Metafont: The Program cited below.

4. TUGboat: ftp://ftp.math.utah.edu/pub/tex/bib/tugboat.bib.

5. TEX and computer typesetting in general:

ftp://ftp.math.utah.edu/pub/tex/bib/texbook1.bib.

6. [Bil87] Neenie Billawala. Write-white printing engines and tuning fonts with Metafont. TUGboat, 8(1):29–32, April 1987.

7. [GMS94] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The LaTEX Companion. Addison-Wesley, Reading, MA, USA, 1994.

8. [Hob89] John D. Hobby. A Metafont-like system with PS output. TUGboat, 10(4):505– 512, December 1989.

9. [Hob92] John D. Hobby. A User’s Manual for MetaPost. Technical Report CSTR-162, AT&T Bell Laboratories, 1992.

10. [Hob93] John D. Hobby. Drawing Graphs with MetaPost. Technical Report CSTR-164, AT&T Bell Laboratories, 1993.

11. [HS91] Samuel P. Harbison and Guy L. Steele Jr. C—A Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1991. An authoritative reference to the C programming language, and a good companion to Kernighan and Ritchie.

12. [KL93] Donald E. Knuth and Silvio Levy. The CWEB System of Structured Documentation, Version 3.0. Addison-Wesley, Reading, MA, USA, 1993.

13. [Knu84] Donald E. Knuth. A torture test for TEX. Report No. STAN-CS-84-1027, Stanford University, Department of Computer Science, 1984.

14. [Knu86a] Donald E. Knuth. A Torture Test for METAFONT. Report No. STAN-CS-86-1095, Stanford University, Department of Computer Science, 1986.

15. [Knu86b] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986.

16. [Knu86c] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986.

17. [Knu86d] Donald E. Knuth. The METAFONTbook, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986.

18. [Knu86e] Donald E. Knuth. METAFONT: The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986.

19. [Knu86f] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986.

20. [Knu89] Donald E. Knuth. The errors of TEX. Software—Practice and Experience, 19(7):607–681, July 1989. This is an updated version of iteKnuth:1988:ET.

21. [Knu90] Donald Knuth. Virtual Fonts: More Fun for Grand Wizards. TUGboat, 11(1):13–23, April 1990.

22. [Knu92] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992.

23. [Lam94] Leslie Lamport. LaTEX: A Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1994. Reprinted with corrections in 1996.

24. [Lia83] Franklin Mark Liang. Word hy-phen-a-tion by com-pu-ter. Technical Report STAN-CS-83-977, Stanford University, August 1983.

25. [Mac91] Pierre A. MacKay. Looking at the pixels: Quality control for 300 dpi laser printer fonts, especially Metafonts. In Robert A. Morris and Jacques Andre, editors, Raster Imaging and Digital Typography II—Papers from the second RIDT meeting, held in Boston, Oct. 14–16, 1991, pages 205–215, New York, 1991. Cambridge University Press.

26. [Spi89] Michael D. Spivak. LAMSTEX, The Synthesis. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1989.

27. [Spi90] Michael D. Spivak. The Joy of TEX—A Gourmet Guide to Typesetting with the AMSTEX macro package. American Mathematical Society, Providence, RI, USA, 2nd revised edition, 1990.

Index 57

Index

#

‘#define’ options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

$

$ expansion in filenames. . . . . . . . . . . . . . . . . . . . . . 13

%

%& magic number . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

-

- starting a filename . . . . . . . . . . . . . . . . . . . . . . . . . . 7

- starts option names . . . . . . . . . . . . . . . . . . . . . . . . . 7 -- starts option names . . . . . . . . . . . . . . . . . . . . . . . . 7

--disable-dump-share configure option. . . . . . 11
--enable-auto-core configure option . . . . . . . . 10

--enable-ipc configure option . . . . . . . . . . . . . . 15 --help common option . . . . . . . . . . . . . . . . . . . . . . . . 7 --verbose common option . . . . . . . . . . . . . . . . . . . . 7 --version common option . . . . . . . . . . . . . . . . . . . . 7 --with-editor=cmd . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --with-epsfwin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-hp2627win . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-mftalkwin . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-regiswin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-suntoolswin . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-tektronixwin. . . . . . . . . . . . . . . . . . . . . . . . 25 --with-unitermwin . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-x. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-x-toolkit=kit . . . . . . . . . . . . . . . . . . . . . . 25 --with-x11 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --with-x11win . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 -base=base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -base=dumpname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -change=chfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 -charcode-format=type . . . . . . . . . . . . . . . . . . 50, 52 -D compiler options . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -dpi=real . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

-file-line-error-style . . . . . . . . . . . . . . . . . . . . . 8

-fmt=dumpname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -fmt=fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -geometry, supported with Xt . . . . . . . . . . . . . . . . 25 -images. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 -ini. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9 -interaction=string . . . . . . . . . . . . . . . . . . . . . . . . 8 -ipc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -ipc-start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -jobname=string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

-kpathsea-debug=number . . . . . . . . . . . . . . . . . . . . . 7

-length=number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -loose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 -lowercase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

-magnification=integer . . . . . . . . . . . . . . . . . 41, 42

-max-pages=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 -mem=dumpname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -mem=mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -min-crossrefs=n . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -mixedcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -mktex=filetype . . . . . . . . . . . . . . . . . . . . . . . . 15, 23 -mltex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -mnemonics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 -no-mktex=filetype . . . . . . . . . . . . . . . . . . . . . 15, 23 -oem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

-output-comment=string . . . . . . . . . . . . . . . . . . . . 15

-output-level=n . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

-overflow-label-offset=points . . . . . . . . . . . . . 26

-page-start=page-spec . . . . . . . . . . . . . . . . . . 41, 42 -parse-first-line . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -progname=string . . . . . . . . . . . . . . . . . . . . . . . . 8, 11 -recorder. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -shell-escape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -show-opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 -strict. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 -style=mftfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 -T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 -terse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 -tex=texprogram . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

-translate-file=tcxfile . . . . . . . . . . . . . . . . . . . . 8

-troff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 32, 34 -underline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 -uppercase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

.

., used for output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 .2602gf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 .aux cross-reference files. . . . . . . . . . . . . . . . . . . . . . 36 .base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 .bbl bibliography files . . . . . . . . . . . . . . . . . . . . . . . 36 .bib bibliography databases . . . . . . . . . . . . . . . . . . 36 .blg BibTEX log file . . . . . . . . . . . . . . . . . . . . . . . . . 36 .fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 .mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 .mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 .mp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 .nnn PostScript figures. . . . . . . . . . . . . . . . . . . . . . . 29 .nnngf generic fonts . . . . . . . . . . . . . . . . . . . . . . . . . 22 .tcx character translation files . . . . . . . . . . . . . 8, 19 .tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 .tfm output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22, 29 .Xdefaults . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 .Xresources . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

^

^^notation, avoiding. . . . . . . . . . . . . . . . . . . . . . . . . 20

\

\bibliography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 \bibliographystyle . . . . . . . . . . . . . . . . . . . . . . . . . 36 \charsubdef and MLTEX . . . . . . . . . . . . . . . . . . . . 18 \countn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 \font and dynamic generation . . . . . . . . . . . . . . . . 14 \fontdimen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \immediate\write18 . . . . . . . . . . . . . . . . . . . . . . . . . 15 \input filenames. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 \mag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 \openout and security. . . . . . . . . . . . . . . . . . . . . . . . 14 \string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 \tracingcharsubdef and MLTEX . . . . . . . . . . . . . 19 \tracinglostchars and MLTEX . . . . . . . . . . . . . . 19 \write18 shell escape extension . . . . . . . . . . . . . . . 15

~

~ expansion in filenames. . . . . . . . . . . . . . . . . . . . . . 13

2

2602gf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

8

8-bit characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

A

abbrv.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 accented character . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 accents, hyphenating words with . . . . . . . . . . . . . . 18 acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 acm.bst. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Ada, WEB for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 additional Make targets . . . . . . . . . . . . . . . . . . . . . . . 4 AFM to TFM conversion . . . . . . . . . . . . . . . . . . . . . 53 afm2tfm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 afmtopl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 aliases for fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 alpha.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 American Mathematical Society, typesetting

system. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 AMSLaTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 AMSTEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 apalike.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Arabic typesetting . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 architecture dependencies . . . . . . . . . . . . . . . . . . . . 11 array limit, fixed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 array sizes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 assembly language routines . . . . . . . . . . . . . . . . . . . . 4 Awk, WEB for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

B

base file, determining . . . . . . . . . . . . . . . . . . . . . . . . 11 base files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 base files, need mode definitions . . . . . . . . . . . . . . 24 base files, plain only . . . . . . . . . . . . . . . . . . . . . . . . . 23 base files, sharing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 bases Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 basic BibTEX style files. . . . . . . . . . . . . . . . . . . . . . . 37 basic fonts and macros . . . . . . . . . . . . . . . . . . . . . . . . 2 batch languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 BDF and GF conversion. . . . . . . . . . . . . . . . . . . . . . 53 beginfig. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 BeginPath ditroff command . . . . . . . . . . . . . . . . . . 33 Berry, Karl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 BIBINPUTS, search path for bib files . . . . . . . . . . . 36 bibliographies, creating. . . . . . . . . . . . . . . . . . . . . . . 36 bibliography. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 bibliography items, cross-referenced . . . . . . . . . . . 36 bibtex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 BibTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 BibTEX collection. . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 BibTEX style files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 BigEndian machines . . . . . . . . . . . . . . . . . . . . . . . . . 11 binaries, linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 blank lines, in TCX files. . . . . . . . . . . . . . . . . . . . . . 20 Bourne shell commands in TEX . . . . . . . . . . . . . . . 15 boxes, memory for . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 breakpoints, memory for . . . . . . . . . . . . . . . . . . . . . . 5 Breitenlohner, Peter . . . . . . . . . . . . . . . . . . . . . . . . . . 1 BSTINPUTS, search path for bst files. . . . . . . . . . . . 36 btex and label extraction. . . . . . . . . . . . . . . . . . . . . 34 btex for MetaPost labels . . . . . . . . . . . . . . . . . . . . . 29 btxdoc.bib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 btxdoc.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 btxhak.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 byte position . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 49 byte swapping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

C

c-sources Makefile target . . . . . . . . . . . . . . . . . . . . . 4 captions, extracting from MetaPost input . . . . . . 34 captions, for MetaPost . . . . . . . . . . . . . . . . . . . . . . . 31 change files, and MFT . . . . . . . . . . . . . . . . . . . . . . . 27 change files, and Tangle . . . . . . . . . . . . . . . . . . . . . . 38 change files, and Weave . . . . . . . . . . . . . . . . . . . . . . 39 changing error messages style . . . . . . . . . . . . . . . . . . 8 character codes, in GFtype output . . . . . . . . . . . . 49 character codes, in PKtype output . . . . . . . . . . . . 47 character codes, in TCX files . . . . . . . . . . . . . . . . . 20 character proofs of fonts . . . . . . . . . . . . . . . . . . . . . . 26 CHARACTER property . . . . . . . . . . . . . . . . . . . . . . . . . . 51 character translation files . . . . . . . . . . . . . . . . . . . . . 19 CHARDP property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHARHT property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHARIC property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 CHARWD property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 chtopx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 class name for Metafont . . . . . . . . . . . . . . . . . . . . . . 25 cm.base. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 cmbase.mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 cmbase.mft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 cmmf.base not recommended . . . . . . . . . . . . . . . . . 23 comments, in TCX files . . . . . . . . . . . . . . . . . . . . . . 20 comments, MFT control . . . . . . . . . . . . . . . . . . . . . . 27 common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 commonalities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 comparing file modification times . . . . . . . . . . . . . 34 compilation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 compile-time options. . . . . . . . . . . . . . . . . . . . . . . . 2, 3 Computer Modern fonts, and Troff . . . . . . . . . . . . 30 Computer Modern macros . . . . . . . . . . . . . . . . . . . . 23 Computer Modern Typefaces, production of … 28 configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 configuration file reading . . . . . . . . . . . . . . . . . . . . . . 9 configuration file values . . . . . . . . . . . . . . . . . . . . . . . 5 configuration, compile-time . . . . . . . . . . . . . . . . . . . . 2

configure --with/--enable options . . . . . . . . . . . 2

CONTENTS.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 control sequence names, space for . . . . . . . . . . . . . . 6 conventions for options, . . . . . . . . . . . . . . . . . . . . . . . 7 conversion, DVI to plain text . . . . . . . . . . . . . . . . . 41 conversion, GF to PK . . . . . . . . . . . . . . . . . . . . . . . . 46 conversion, GF to plain text . . . . . . . . . . . . . . . . . . 48 conversion, PK to GF . . . . . . . . . . . . . . . . . . . . . . . . 46 conversion, PK to plain text . . . . . . . . . . . . . . . . . . 47 conversion, property list to TFM . . . . . . . . . . . . . . 52 conversion, property list to VF. . . . . . . . . . . . . . . . 53 conversion, TFM to property list . . . . . . . . . . . . . . 50 conversion, VF to VPL . . . . . . . . . . . . . . . . . . . . . . . 52 copyright notices . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 core dump from filename. . . . . . . . . . . . . . . . . . . . . 10 Cork encoding and ISO input . . . . . . . . . . . . . . . . . 20 creating memory dumps . . . . . . . . . . . . . . . . . . . . . . 10 cross-referenced bibliography items . . . . . . . . . . . . 36 cross-references, omitting . . . . . . . . . . . . . . . . . . . . . 39 CTRL-\ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 current directory, used for output . . . . . . . . . . . . . . 9 Curtis, Pavel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Cweb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 CWEB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

D

D c ditroff graphics . . . . . . . . . . . . . . . . . . . . . . . . . . 33 date and time, in memory dumps . . . . . . . . . . . . . 12 debugging DVI utilities. . . . . . . . . . . . . . . . . . . . . . . 42 debugging flags, specifying. . . . . . . . . . . . . . . . . . . . . 7 decimal character codes, in TCX files. . . . . . . . . . 20 dependencies, hardware . . . . . . . . . . . . . . . . . . . . . . 11 design-size units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 device definitions, for Metafont . . . . . . . . . . . . . . . 24 device-independent width . . . . . . . . . . . . . . . . . 47, 49 directory structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 DISPLAY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 ditroff output, converting to MPX . . . . . . . . . . . . 33 DMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 33 DMP, invoked by MakeMPX . . . . . . . . . . . . . . . . . 32 dmp.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 dot files, written by TEX programs . . . . . . . . . . . . 14 downloading of fonts for MetaPost labels . . . . . . 30 dpost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 DrawingServant . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 DrawPath ditroff command. . . . . . . . . . . . . . . . . . . . 33 dump file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 dumping memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 DVI comment, specifying . . . . . . . . . . . . . . . . . . . . . 15 DVI files, converting to MPX . . . . . . . . . . . . . . . . . 33 DVI files, creating multiple . . . . . . . . . . . . . . . . . . . 21 DVI files, explained . . . . . . . . . . . . . . . . . . . . . . . . . . 45 DVI format definition . . . . . . . . . . . . . . . . . . . . . . . . 41 DVI opcodes, showing. . . . . . . . . . . . . . . . . . . . . . . . 42 DVI utilities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 dvicopy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 dvitomp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 DVITOMP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 DVItoMP, invoked by MakeMPX . . . . . . . . . . . . . 32 dvitype DVI validation . . . . . . . . . . . . . . . . . . . . . . 41 dvitype output example . . . . . . . . . . . . . . . . . . . . . . 42 dvitype.web . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 dx horizontal escapement . . . . . . . . . . . . . . . . . 47, 49 dy vertical escapement . . . . . . . . . . . . . . . . . . . . . . . 49 dynamic array allocation . . . . . . . . . . . . . . . . . . . . . . 6 dynamic Metafont mode definitions with smode

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 dynamic packing variable. . . . . . . . . . . . . . . . . . . . . 47

E

e response at error prompt . . . . . . . . . . . . . . . . . . . 12 e-circumflex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 e-TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 e.mft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 EC fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 22 editing of bitmap fonts . . . . . . . . . . . . . . . . . . . . . . . 53 editor invoked at error . . . . . . . . . . . . . . . . . . . . . . . 12 eight-bit characters in filenames. . . . . . . . . . . . . . . 13 endian dependencies . . . . . . . . . . . . . . . . . . . . . . . . . 11 eoc GF command. . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Eplain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 epsf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 errors, editor invoked at . . . . . . . . . . . . . . . . . . . . . . 12 escapement, horizontal. . . . . . . . . . . . . . . . . . . . 47, 49 escapement, vertical . . . . . . . . . . . . . . . . . . . . . . . . . 49 etex and label extraction. . . . . . . . . . . . . . . . . . . . . 34 etex for MetaPost labels . . . . . . . . . . . . . . . . . . . . . 29 executables, shared initial and virgin . . . . . . . . . . . 9 expanded plain format . . . . . . . . . . . . . . . . . . . . . . . 18 extensions to TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 extra_mem_bot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

FACE property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 FAMILY property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Ferguson, Michael . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 file formats for fonts . . . . . . . . . . . . . . . . . . . . . . . . . 45 file mtimes, comparing . . . . . . . . . . . . . . . . . . . . . . . 34 file recorder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 File-handling TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 filename conventions, in input files . . . . . . . . . . . . 13 filenames starting with ‘-’ . . . . . . . . . . . . . . . . . . . . . 7 first line of the main input file . . . . . . . . . . . . . . . . . 8 fixed-point arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . 4 FIXPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 flag byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 floating-point arithmetic. . . . . . . . . . . . . . . . . . . . . . . 4 floating-point values . . . . . . . . . . . . . . . . . . . . . . . . . 12 fmt file, determining . . . . . . . . . . . . . . . . . . . . . . . . . 11 fmt files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 fmt files, sharing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 fmts Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 font aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 font character code, translating . . . . . . . . . . . . . . . 20 font design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 font downloading for MetaPost labels. . . . . . . . . . 30 font file formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 font proofs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 font utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 font utilities, non-Web2c . . . . . . . . . . . . . . . . . . . . . 53 font_mem_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 fontinst, for creating virtual fonts. . . . . . . . . . . . 53 fonts, basic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 format files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 formats for TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 formats Make target . . . . . . . . . . . . . . . . . . . . . . . . . . 4 fraction routines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Free Software Foundation documentation system

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 freedom of Web2c. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ftp.math.utah.edu . . . . . . . . . . . . . . . . . . . . . . . . . . 36

G

generating source specials . . . . . . . . . . . . . . . . . . . . 16 geometric designs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 geometric font scaling . . . . . . . . . . . . . . . . . . . . . . . . 45 geometry for Metafont . . . . . . . . . . . . . . . . . . . . . . . 25 getopt_long_only . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 GF files, explained . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 GF files, output by Metafont . . . . . . . . . . . . . . . . . 22 GF format definition . . . . . . . . . . . . . . . . . . . . . . . . . 45 GF output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 GF, converting PK to . . . . . . . . . . . . . . . . . . . . . . . . 46 GF, converting to PK . . . . . . . . . . . . . . . . . . . . . . . . 46 gftodvi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 gftopk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 gftopxl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 gftype GF validation . . . . . . . . . . . . . . . . . . . . . . . . 48 gftype.web . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 glue ratio representations . . . . . . . . . . . . . . . . . . . . . 12 glue, memory for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 glyph substitutions . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gray font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Gruff, Billy Goat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gsftopk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

H

                                                HackyInputFileNameForCoreDump.tex. . . . . . . . . 10

Harbison, Samuel P. . . . . . . . . . . . . . . . . . . . . . . . . . 12 hardware and memory dumps. . . . . . . . . . . . . . . . . 11 hash table, increasing size of . . . . . . . . . . . . . . . . . . . 6 hash_extra . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 headerbyte information . . . . . . . . . . . . . . . . . . . . . . 51 height, in pixels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 help, online. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Henry, Patrick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Herberts, Mathias . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 hex character codes, in TCX files . . . . . . . . . . . . . 20 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Hobby, John . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 34 horizontal escapement . . . . . . . . . . . . . . . . . . . . 47, 49 hp2627 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 human languages, supported in TEX. . . . . . . . . . . 18 human-readable text, converting DVI to . . . . . . . 41 human-readable text, converting GF to . . . . . . . . 48 human-readable text, converting PK to . . . . . . . . 47 human-readable text, converting TFM to . . . . . . 50 human-readable text, converting VF to . . . . . . . . 52 hyphenation and languages . . . . . . . . . . . . . . . . . . . 18 hyphenation patterns, creating . . . . . . . . . . . . . . . . 21

I

ice cream. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 identifier case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 identifier collisions . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 identifier length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 identifiers with underlines . . . . . . . . . . . . . . . . . . . . 39 ieeetr.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 il1-t1.tcx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 il2-t1.tcx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Info format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 initial form, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . 8 initial Metafont. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 initial MetaPost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 initial programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 initial TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 initializations, lengthy. . . . . . . . . . . . . . . . . . . . . . . . 10 input filenames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 install-bases Make target . . . . . . . . . . . . . . . . . . . 4 install-fmts Make target . . . . . . . . . . . . . . . . . . . . 4 install-formats Make target . . . . . . . . . . . . . . . . . 4 install-mems Make target . . . . . . . . . . . . . . . . . . . . 4 installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 interaction mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 international characters . . . . . . . . . . . . . . . . . . . . . . 19 introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 IPC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 IPC_DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 21

J

job name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

K

kerning table, in TFM files . . . . . . . . . . . . . . . . . . . 51 keyboard character code, translating . . . . . . . . . . 20 Knuth, Donald E. . . . . . . . . . . . . . . . . . . . . . . . . . 1, 27 KPATHSEA_DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 KRN property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

L

label font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 LABEL property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 labels, extracting from MetaPost input . . . . . . . . 34 labels, for MetaPost. . . . . . . . . . . . . . . . . . . . . . . . . . 31 LAMSTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 language support in TEX . . . . . . . . . . . . . . . . . . . . . 18 languages, hyphenation rules for . . . . . . . . . . . . . . 21 LaTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 left side bearing . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 49 legalisms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 licensing terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 LIG property . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 ligature table, in TFM files . . . . . . . . . . . . . . . . . . . 51 LIGTABLE property . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 linking binaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 links to binaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 literate programming. . . . . . . . . . . . . . . . . . . . . . . . . 38 litprog@shsu.edu . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 LittleEndian machines . . . . . . . . . . . . . . . . . . . . . . . 11 log file, BibTEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

M

machine dependencies . . . . . . . . . . . . . . . . . . . . . . . . 11 machine-readable, converting property lists to

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52, 53 MacKay, Pierre . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 macro packages, major TEX . . . . . . . . . . . . . . . . . . 17 macros, basic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 macros, predefining in memory dumps . . . . . . . . . 10 magnification . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 main_memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Make targets, additional. . . . . . . . . . . . . . . . . . . . . . . 4 makempx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 MAKEMPX_BINDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Martin, Rick . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Mathematical Reviews . . . . . . . . . . . . . . . . . . . . . . . 17 mathematical typesetting. . . . . . . . . . . . . . . . . . . . . 14 mem file, determining . . . . . . . . . . . . . . . . . . . . . . . . 11 mem files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 mem files, plain only . . . . . . . . . . . . . . . . . . . . . . . . . 31 mem files, sharing . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 memory dump to use, determining . . . . . . . . . . . . 11 memory dumps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 memory dumps and hardware. . . . . . . . . . . . . . . . . 11 memory dumps, contain date and time . . . . . . . . 12 memory dumps, creating . . . . . . . . . . . . . . . . . . . . . 10 mems Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 meta characters in filenames . . . . . . . . . . . . . . . . . . 13 Metafont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Metafont geometry. . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Metafont graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Metafont input files . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Metafont invocation. . . . . . . . . . . . . . . . . . . . . . . . . . 22 Metafont meets PostScript . . . . . . . . . . . . . . . . . . . 29 Metafont online support, new devices. . . . . . . . . . 26 Metafont source, prettyprinting . . . . . . . . . . . . . . . 27 Metafont, compatibility in MetaPost . . . . . . . . . . 31 Metafont, initial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Metafont, MetaPost, and TEX . . . . . . . . . . . . . . . . . 9 Metafont, virgin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 MetaPost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 MetaPost and plain Metafont compatibility …. 31 MetaPost input files . . . . . . . . . . . . . . . . . . . . . . . . . 29 MetaPost input, extracting labels from . . . . . . . . 34 MetaPost invocation . . . . . . . . . . . . . . . . . . . . . . . . . 29 MetaPost labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 MetaPost source, prettyprinting. . . . . . . . . . . . . . . 28 MetaPost, initial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 MetaPost, TEX, and Metafont . . . . . . . . . . . . . . . . . 9 MetaPost, virgin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 mf.base. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 MFEDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 mfplain.mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 mfput . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 mft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 mftalk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 MFTERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 mftmac.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 mktexmf, disabling . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 mktextfM, disabling . . . . . . . . . . . . . . . . . . . . . . . . . . 14 mltex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 MLTEX, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 mode needed to run Metafont. . . . . . . . . . . . . . . . . 22 mode_def. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 mode_setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 modes file needed for Metafont. . . . . . . . . . . . . . . . 24 modes.mf recommended modes file . . . . . . . . . . . . 24 Morgan, Tim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Morris, Bob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 MPEDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 mpgraph.ps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mpman.ps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mpost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mpost, reason for name change . . . . . . . . . . . . . . . . . 2 mpost.mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 mpout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mproof.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 MPSUPPORT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 MPto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 MPto, invoked by MakeMPX . . . . . . . . . . . . . . . . . 31 MPTOTEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 MPTOTR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mptrap Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 mptrap test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 mptrap.readme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 mpx file, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 MPX files, converting from DVI files . . . . . . . . . . 33 MPX files, creating from ditroff output . . . . . . . . 33 mpxerr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mpxerr.dvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mpxerr.log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mpxerr.t. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mpxerr.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 mtimes of files, comparing . . . . . . . . . . . . . . . . . . . . 34 Multi-lingual TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 multiple DVI files, creating . . . . . . . . . . . . . . . . . . . 21

N

N tilde . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 new graphics support for Metafont . . . . . . . . . . . . 26 NEWER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 newer file comparison . . . . . . . . . . . . . . . . . . . . . . . . 34 newrow GF command . . . . . . . . . . . . . . . . . . . . . . . . 49 next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 NO_X11WIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 non-Unix system, compiling on . . . . . . . . . . . . . . . . 4 NUL, not allowed in filenames . . . . . . . . . . . . . . . . 13

O

octal character codes, in TCX files . . . . . . . . . . . . 20 offset for overflow labels . . . . . . . . . . . . . . . . . . . . . . 26 older-than, file comparisons. . . . . . . . . . . . . . . . . . . 34 Omega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 online Metafont graphics . . . . . . . . . . . . . . . . . . . . . 25 opcodes, showing DVI. . . . . . . . . . . . . . . . . . . . . . . . 42 optical font scaling. . . . . . . . . . . . . . . . . . . . . . . . . . . 45 option conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 origin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 output file location. . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 output files, written by TEX programs . . . . . . . . . 14 output_comment for DVI files . . . . . . . . . . . . . . . . . 15 overflow label offset . . . . . . . . . . . . . . . . . . . . . . . . . . 26

P

packet length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 page, starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 parsing the first line. . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Pascal, creating from WEB . . . . . . . . . . . . . . . . . . . 38 patgen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 path searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 path searching debugging. . . . . . . . . . . . . . . . . . . . . . 7 PDF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 pdfTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 permissions, legal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 PFA and PFB conversion. . . . . . . . . . . . . . . . . . . . . 53 PiCTEX, increasing memory for . . . . . . . . . . . . . . . . 5 picture expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 pixel height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 pixel width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 PK bitmaps from PostScript. . . . . . . . . . . . . . . . . . 53 PK files, explained . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PK files, not output by Metafont. . . . . . . . . . . . . . 22 PK format definition . . . . . . . . . . . . . . . . . . . . . . . . . 45 PK, converting GF to . . . . . . . . . . . . . . . . . . . . . . . . 46 PK, converting to GF . . . . . . . . . . . . . . . . . . . . . . . . 46 pktogf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 pktype PK validation . . . . . . . . . . . . . . . . . . . . . . . . 47 pktype.web . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 PL files, explained . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 plain Metafont, compatibility in MetaPost . . . . . 31 plain text, converting DVI to . . . . . . . . . . . . . . . . . 41 plain text, converting GF to . . . . . . . . . . . . . . . . . . 48 plain text, converting PK to . . . . . . . . . . . . . . . . . . 47 plain text, converting TFM to . . . . . . . . . . . . . . . . 50 plain text, converting VF to . . . . . . . . . . . . . . . . . . 52 plain.base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 plain.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 plain.fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 plain.mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 plain.mft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 pltotf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 pool file, writing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Poole, Simon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 pooltype. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 PostScript fonts, and Troff. . . . . . . . . . . . . . . . . . . . 30 PostScript meets Metafont . . . . . . . . . . . . . . . . . . . 29 PostScript output. . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 PostScript to PK bitmaps . . . . . . . . . . . . . . . . . . . . 53 PostScript Type 1 font conversion . . . . . . . . . . . . . 53 PostScript, and font scaling. . . . . . . . . . . . . . . . . . . 45 predefined macros and memory dumps . . . . . . . . 10 preloaded executables, creating . . . . . . . . . . . . . . . 10 preloaded executables, not recommended . . . . . . 10 prettyprinting Metafont source. . . . . . . . . . . . . . . . 27 prettyprinting WEB programs . . . . . . . . . . . . . . . . 39 primitives, new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 printable characters, specifying . . . . . . . . . . . . . . . 20 printer characteristics, for Metafont . . . . . . . . . . . 24 production use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 program name, determines memory dump . . . . . 11 program names, special . . . . . . . . . . . . . . . . . . . . 8, 15 prologues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 prologues and Troff in MetaPost . . . . . . . . . . . . . 32 prologues, and EPSF output . . . . . . . . . . . . . . . . . 30 proof mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 proof sheets, of fonts . . . . . . . . . . . . . . . . . . . . . . . . . 26 property list format . . . . . . . . . . . . . . . . . . . . . . . . . . 50 property list, converting TFM to . . . . . . . . . . . . . . 50 property list, converting VF to virtual . . . . . . . . . 52 ps2pk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 psfonts.map, read by MetaPost. . . . . . . . . . . . . . . 30 PXL files, explained. . . . . . . . . . . . . . . . . . . . . . . . . . 45 pxtoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

Q

quit character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

R

Raichle, Bernd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 reading, additional . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 readonly directory, running TEX in . . . . . . . . . . . . . 9 reallocation of arrays . . . . . . . . . . . . . . . . . . . . . . . . . . 6 redefined character substitutions . . . . . . . . . . . . . . 19 reference pixel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 regis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Regis graphics support . . . . . . . . . . . . . . . . . . . . . . . 25 regression testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 repeated rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 representation of strings . . . . . . . . . . . . . . . . . . . . . . 40 right side bearing. . . . . . . . . . . . . . . . . . . . . . . . . 47, 49 right-to-left typesetting . . . . . . . . . . . . . . . . . . . . . . 21 Rokicki, Tomas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 run length encoded bitmaps . . . . . . . . . . . . . . . 48, 49 runtime options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

scaled pixels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 49 scaling of fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 scanned images of fonts . . . . . . . . . . . . . . . . . . . . . . 53 security, and \openout . . . . . . . . . . . . . . . . . . . . . . . 14 security, and output files . . . . . . . . . . . . . . . . . . . . . 14 security, and shell escapes . . . . . . . . . . . . . . . . . . . . 15 SetColor ditroff command. . . . . . . . . . . . . . . . . . . . 33 shapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 sharing executables, and preloading . . . . . . . . . . . 10 sharing memory dumps . . . . . . . . . . . . . . . . . . . . . . 11 shell commands in TEX . . . . . . . . . . . . . . . . . . . . . . 15 shell_escape enabling in TEX. . . . . . . . . . . . . . . . 15 siam.bst. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 side bearings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 49 SIGQUIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 slant font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 slides, producing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 SliTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 small Metafont memory and modes . . . . . . . . . . . 24 smode and dynamic Metafont mode definition . . 24 sockets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 space-terminated filenames . . . . . . . . . . . . . . . . . . . 13 Spiderweb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Stallman, Richard . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 starting page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 42 Steele Jr., Guy L. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 strategy, overall. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 string numbers, displaying . . . . . . . . . . . . . . . . . . . . 39 string pool, writing . . . . . . . . . . . . . . . . . . . . . . . . . . 38 string representation . . . . . . . . . . . . . . . . . . . . . . . . . 40 style design, for BibTEX . . . . . . . . . . . . . . . . . . . . . . 36 style files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 substitutions of font glyphs . . . . . . . . . . . . . . . . . . . 18 sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 sun-gfx.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Suntools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 SunView . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 swap space, as array limit . . . . . . . . . . . . . . . . . . . . . 6 swapping bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 syntax of TCX files . . . . . . . . . . . . . . . . . . . . . . . . . . 20 system C library function . . . . . . . . . . . . . . . . . . . . 15

T

T1 encoding and ISO input . . . . . . . . . . . . . . . . . . . 20 Tachikawa, Elizabeth. . . . . . . . . . . . . . . . . . . . . . . . . 46 tangle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 targets, additional Make. . . . . . . . . . . . . . . . . . . . . . . 4 TCX character translation files . . . . . . . . . . . . . . . . . 19 tek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Tektronix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Tektronix 4014 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 terminator for filenames . . . . . . . . . . . . . . . . . . . . . . 13 terse output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 TeX--XeT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 tex.fmt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 TEX, bibliographies for . . . . . . . . . . . . . . . . . . . . . . . 36 TEX, creating from Metafont . . . . . . . . . . . . . . . . . 27 TEX, creating from WEB . . . . . . . . . . . . . . . . . . . . . 39 TEX, description of . . . . . . . . . . . . . . . . . . . . . . . . . . 14 TEX, extensions to . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 TEX, format packages for . . . . . . . . . . . . . . . . . . . . . 17 TEX, initial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 TEX, input files found . . . . . . . . . . . . . . . . . . . . . . . . 14 TEX, invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 TEX, Metafont, and MetaPost . . . . . . . . . . . . . . . . . 9 TEX, virgin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 TEX, Web2c implementation of . . . . . . . . . . . . . . . . 1 TEXBIB, search path for bib files. . . . . . . . . . . . . . . 36 TEXEDIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 texfonts.map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Texinfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 texmf.cnf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 texmfmp.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

texmfmp.c and openoutnameok. . . . . . . . . . . . . . . . 14

TEXMFOUTPUT, used if ‘.’ unwritable. . . . . . . . . . . . . 9 texput . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 text, extracting from MetaPost input. . . . . . . . . . 34 TFM files, converting property lists to. . . . . . . . . 52 TFM files, explained . . . . . . . . . . . . . . . . . . . . . . . . . 45 TFM files, memory for . . . . . . . . . . . . . . . . . . . . . . . . 5 TFM files, output by Metafont. . . . . . . . . . . . . . . . 22 TFM files, output by MetaPost . . . . . . . . . . . . . . . 29 TFM width of characters . . . . . . . . . . . . . . . . . 47, 49 tftopl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 three programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 time and date, in memory dumps . . . . . . . . . . . . . 12 title font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 toolkits, X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 torture tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 translation file for TEX, specifying. . . . . . . . . . . . . . 8 translation from WEB to C. . . . . . . . . . . . . . . . . . . . 1 trap Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 trap test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 trapman.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 trchars.adj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 34 Trickey, Howard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 trip Make target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 trip test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 tripman.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 triptrap Make target. . . . . . . . . . . . . . . . . . . . . . . . . 4 TROFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Troff, and MetaPost . . . . . . . . . . . . . . . . . . . . . . . . . 30 Troff, WEB for . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 Trojan horses and TEX programs. . . . . . . . . . . . . . 14 TUGboat bibliography . . . . . . . . . . . . . . . . . . . . . . . 36 Type 1 conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 type design, personal. . . . . . . . . . . . . . . . . . . . . . . . . 24 type programs, DVI. . . . . . . . . . . . . . . . . . . . . . . . . . 41 type programs, GF . . . . . . . . . . . . . . . . . . . . . . . . . . 48 type programs, PK . . . . . . . . . . . . . . . . . . . . . . . . . . 47 type programs, pool. . . . . . . . . . . . . . . . . . . . . . . . . . 39 typeface families . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 typeface specimen sheets . . . . . . . . . . . . . . . . . . . . . 53 typesetting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

U

undump . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Unicode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 uniterm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 unsrt.bst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 using local codepage to display messages. . . . . . . . 8

V

validation, of DVI files . . . . . . . . . . . . . . . . . . . . . . . 41 validation, of GF files . . . . . . . . . . . . . . . . . . . . . . . . 48 validation, of PK files . . . . . . . . . . . . . . . . . . . . . . . . 47 validation, of TFM files . . . . . . . . . . . . . . . . . . . . . . 50 validation, of VF files . . . . . . . . . . . . . . . . . . . . . . . . 52 verbatimtex MetaPost command . . . . . . . . . . . . . 34 verbose BibTEX output, suppressing. . . . . . . . . . . 36 verbosity, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 version number, finding . . . . . . . . . . . . . . . . . . . . . . . 7 vertical escapement . . . . . . . . . . . . . . . . . . . . . . . . . . 49 VF files, converting property lists to. . . . . . . . . . . 53 vftovp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 virgin Metafont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 virgin MetaPost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 virgin programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 virgin TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 virtual font creation. . . . . . . . . . . . . . . . . . . . . . . . . . 53 virtual fonts, expanding . . . . . . . . . . . . . . . . . . . . . . 41 virtual-fonts.knuth. . . . . . . . . . . . . . . . . . . . . . . . 45 virtualfonts.txt . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 vptovf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

W

weave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 WEB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 WEB pool files, displaying. . . . . . . . . . . . . . . . . . . . 39 WEB programs, compiling. . . . . . . . . . . . . . . . . . . . 38 WEB programs, typesetting . . . . . . . . . . . . . . . . . . 39 WEB2C, search path for TCX files . . . . . . . . . . . . . . 20 Weber, Olaf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 webmac.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 webman.tex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 whitespace, in TCX files. . . . . . . . . . . . . . . . . . . . . . 20 whitespace-terminated filenames . . . . . . . . . . . . . . 13 width, device-independent. . . . . . . . . . . . . . . . . 47, 49 width, in pixels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 word processor, not . . . . . . . . . . . . . . . . . . . . . . . . . . 14 writing memory dumps. . . . . . . . . . . . . . . . . . . . . . . 10

X

X bitmap fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 X class name for Metafont. . . . . . . . . . . . . . . . . . . . 25 x offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 X resources. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 X toolkits and Metafont . . . . . . . . . . . . . . . . . . . . . . 25 x X ditroff device control . . . . . . . . . . . . . . . . . . . . . 33 xampl.bib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 xbfe, bitmap font editor. . . . . . . . . . . . . . . . . . . . . . 53 xfed, bitmap font editor. . . . . . . . . . . . . . . . . . . . . . 53 xfedor, bitmap font editor . . . . . . . . . . . . . . . . . . . 53 Xlib . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Xlib support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Xt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Xt support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 xterm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Y

y offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 configure options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Compile-time options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Additional targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.4 Trip, trap, and mptrap: Torture tests . . . . . . . . . . . . . . . . . . . . . 4 2.5 Runtime options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Commonalities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.1 Option conventions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Path searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.4 Output file location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.5 Three programs: Metafont, MetaPost, and TEX. . . . . . . . . . . . 9

3.5.1 Initial and virgin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3.5.1.1 Preloaded executables . . . . . . . . . . . . . . . . . 10 3.5.2 Memory dumps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.5.2.1 Creating memory dumps . . . . . . . . . . . . . . 10 3.5.2.2 Determining the memory dump to use . . 11 3.5.2.3 Hardware and memory dumps . . . . . . . . . 11

3.5.3 Editor invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.5.4 \input filenames . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4 TEX: Typesetting . . . . . . . . . . . . . . . . . . . . . . . . . 14

4.1 tex invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 initex invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.3 virtex invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.4 Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.5 Languages and hyphenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

4.5.1 MLTEX: Multi-lingual TEX . . . . . . . . . . . . . . . . . . . . . 18 4.5.1.1 \charsubdef: Character substitutions . . 18 4.5.1.2 \tracingcharsubdef: Substitution

diagnostics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4.5.2 TCX files: Character translations . . . . . . . . . . . . . . . 19 4.5.3 Patgen: Creating hyphenation patterns. . . . . . . . . . 20

4.6 IPC and TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.7 TEX extensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

5 Metafont: Creating typeface families . . . . . . . 22

5.1 mf invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 5.2 inimf invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.3 virmf invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 5.4 Modes: Device definitions for Metafont . . . . . . . . . . . . . . . . . . 24 5.5 Online Metafont graphics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.6 GFtoDVI: Character proofs of fonts . . . . . . . . . . . . . . . . . . . . . 26 5.7 MFT: Prettyprinting Metafont source. . . . . . . . . . . . . . . . . . . . 26

6 MetaPost: Creating technical illustrations . . 29

6.1 mpost invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.2 inimpost invocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 6.3 virmpost invocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 6.4 MakeMPX: Support MetaPost labels . . . . . . . . . . . . . . . . . . . . 31 6.5 DVItoMP: DVI to MPX conversion . . . . . . . . . . . . . . . . . . . . . . 32 6.6 DMP: Ditroff to MPX conversion. . . . . . . . . . . . . . . . . . . . . . . . 33 6.7 MPto: Extract labels from MetaPost input . . . . . . . . . . . . . . . 34 6.8 Newer: Compare file modification times . . . . . . . . . . . . . . . . . . 34

7 BibTEX: Bibliographies . . . . . . . . . . . . . . . . . . . 36

7.1 BibTEX invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.2 Basic BibTEX style files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

8 WEB: Literate programming . . . . . . . . . . . . . . 38

8.1 Tangle: Translate WEB to Pascal . . . . . . . . . . . . . . . . . . . . . . . 38 8.2 Weave: Translate WEB to TEX. . . . . . . . . . . . . . . . . . . . . . . . . . 39 8.3 Pooltype: Display WEB pool files . . . . . . . . . . . . . . . . . . . . . . . 39

9 DVI utilities. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

9.1 DVIcopy: Canonicalize virtual font references . . . . . . . . . . . . 41 9.2 DVItype: Plain text transliteration of DVI files . . . . . . . . . . . 41

9.2.1 DVItype output example . . . . . . . . . . . . . . . . . . . . . . . 42

10 Font utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

10.1 Font file formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 10.2 GFtoPK: Generic to packed font conversion . . . . . . . . . . . . . 46 10.3 PKtoGF: Packed to generic font conversion . . . . . . . . . . . . . 46 10.4 PKtype: Plain text transliteration of packed fonts . . . . . . . 46 10.5 GFtype: Plain text transliteration of generic fonts . . . . . . . 48 10.6 TFtoPL: TEX font metric to property list conversion . . . . . 49 10.7 PLtoTF: Property list to TEX font metric conversion. . . . . 51 10.8 VFtoVP: Virtual font to virtual property lists . . . . . . . . . . . 52 10.9 VPtoVF: Virtual property lists to virtual font . . . . . . . . . . . 52 10.10 Font utilities available elsewhere. . . . . . . . . . . . . . . . . . . . . . . 53

Appendix A Legalisms . . . . . . . . . . . . . . . . . . . . . 54 Appendix B References . . . . . . . . . . . . . . . . . . . . 55

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57